### v1

# Log-concavity of polynomials arising from equivariant cohomology

Abstract.
We study the equivariant cohomology classes of torus-equivariant subvarieties of the space of matrices. For a large class of torus actions, we prove that the polynomials representing these classes (up to suitably changing signs) are covolume polynomials in the sense of Aluffi. We study the cohomology rings of complex varieties in terms of Macaulay inverse systems over ${\mathbb{Z}}$. As applications, we show that under certain conditions, the Macaulay dual generator is a denormalized Lorentzian polynomial in the sense of Brändén and Huh, and we give a characteristic-free extension (over $\mathbb{Z}$) of the result of Khovanskii and Pukhlikov describing the cohomology ring of toric varieties in terms of volume polynomials.
Key words and phrases:
equivariant cohomology, multidegrees, cohomology rings, Richardson polynomials, Lorentzian polynomials, covolume polynomials, log-concavity, Schubert polynomials, Richardson varieties, Schubert varieties
2020 Mathematics Subject Classification:
14M15, 14C15, 14C17, 13H15, 52B40
Jacob Matherne received support from a Simons Foundation Travel Support for Mathematicians Award MPS-TSM-00007970.
1. Introduction
A sequence of real numbers $a_{0},a_{1},\ldots,a_{n}$ is called log-concave if $a_{i}^{2}\geq a_{i-1}a_{i+1}$ for all $11\leq i\leq n-$ 1. Log-concave sequences naturally appear throughout algebra, combinatorics, and geometry; for thorough references on log-concavity and related concepts, we suggest [Bre89, Sta89, Brä15]. Recently, the theory of Lorentzian polynomials was introduced by Brändén and Huh [BH20] (and in [AGV21, ALOGV24a, ALOGV24b] under the name of completely log-concave polynomials) and they have been instrumental in proving log-concavity and related results throughout mathematics, see, e.g., [BH20, EH20, MS21, HMMSD22, BL23, BLP23, Ros23, ALOGV24b, BES24, HMV24, KMSD24, MMS24, RU24].
The prototypical examples of Lorentzian polynomials are the volume polynomials of projective varieties [BH20]. Likewise, the covolume polynomials of Aluffi [Alu24] are the prototypical examples of the dually Lorentzian polynomials of Ross, Süß, and Wannerer [RSW23]. We introduce a new family of polynomials that specialize to a number of important polynomials in algebraic combinatorics, and we prove that they are covolume polynomials.
For two permutations $u,w\in S_{n}$ with $w\geq u$ in Bruhat order, we define the double Richardson polynomial as the product of double Schubert polynomials
$\mathfrak{R}_{w/u}({\textbf{t}},{\textbf{s}})\;=\;\mathfrak{S}_{u}({\textbf{t}},{\textbf{s}})\mathfrak{S}_{w_{0}w}({\textbf{t}},{\textbf{s}}^{\prime}),$
where ${\textbf{s}}^{\prime}=(s_{n},\ldots,s_{1})$ denotes the reverse of ${\textbf{s}}=(s_{1},\ldots,s_{n})$. The double Richardson polynomial represents the torus-equivariant class of matrix Richardson varieties in the cohomology ring $\text{H}_{T}^{\bullet}({\rm Mat}_{n,n})$ of the space of $n\times n$ matrices with the standard action of the torus $T=({\mathbb{C}}^{*})^{n}\times({\mathbb{C}}^{*})^{n}$. Double Richardson polynomials specialize to (double) Schubert polynomials:
$\mathfrak{S}_{u}({\textbf{t}},{\textbf{s}})\;=\;\mathfrak{R}_{w_{0}/u}({\textbf{t}},{\textbf{s}})\qquad\text{and}\qquad\mathfrak{S}_{u}({\textbf{t}},\mathbf{0})\;=\;\mathfrak{S}_{u}({\textbf{t}})\;=\;\mathfrak{R}_{w_{0}/u}({\textbf{t}},\mathbf{0}).$
The Richardson polynomial $\mathfrak{R}_{w/u}({\textbf{t}})=\mathfrak{R}_{w/u}({\textbf{t}},\mathbf{0})$ is closely related to the skew Schubert polynomial $\mathfrak{S}_{w/u}({\textbf{t}})$ of Lenart and Sottile [LS03] since both polynomials yield the same class in the cohomology ring $\text{H}^{\bullet}({{\mathcal{F}\ell}_{n}})$ of the flag variety ${{\mathcal{F}\ell}_{n}}$. But we point out that $\mathfrak{R}_{w/u}({\textbf{t}})$ and $\mathfrak{S}_{w/u}({\textbf{t}})$ may not be equal in general (see Section missing 5).
Theorem A (Theorem 5.2).
The (sign-changed) double Richardson polynomial $\mathfrak{R}_{w/u}({\textbf{t}},-{\textbf{s}})$ is a covolume polynomial.
Dually Lorentzian polynomials enjoy two nice combinatorial properties (see Section missing 5): their supports are M-convex, and they are discretely log-concave. A homogeneous polynomial $h=\sum_{\mathbf{n}}a_{\mathbf{n}}{\textbf{t}}^{\mathbf{n}}$ of degree $d$ with nonnegative coefficients is said to have M-convex support if $\mathrm{supp}(h)$ is the set of integer points of a generalized permutohedron in the sense of [Pos09], and it is said to be discretely log-concave if $a_{\mathbf{n}}^{2}\geq a_{{\mathbf{n}}+{\mathbf{e}}_{i}-{\mathbf{e}}_{j}}a_{{\mathbf{n}}-{\mathbf{e}}_{i}+{\mathbf{e}}_{j}}$ for all ${\mathbf{n}}\in\mathbb{N}^{n}$ and all $1\leq i,j\leq n$.
[…]
The following polynomials have M-convex support and are discretely log-concave:
  1. (i)
    (sign-changed) Double Richardson polynomials $\mathfrak{R}_{w/u}({\textbf{t}},-{\textbf{s}})$.
  2. (ii)
    Richardson polynomials $\mathfrak{R}_{w/u}({\textbf{t}})$.
  3. (iii)
    (sign-changed) Double Schubert polynomials $\mathfrak{S}_{u}({\textbf{t}},-{\textbf{s}})$.
  4. (iv)
    Schubert polynomials $\mathfrak{S}_{u}({\textbf{t}})$.
We note that the M-convexity in Section missing 1(iii) recovers a result of [CCRMMn23] and both the M-convexity and the discrete log-concavity in Section missing 1(iv) recover results of [HMMSD22]. To the best of our knowledge, the discrete log-concavity of double Schubert polynomials is new.
The proof of Theorem A follows from a general theorem that we prove in Section missing 4. More precisely, we prove the following result regarding the equivariant cohomology classes of torus-equivariant subvarieties of the space of matrices.
Theorem B (Theorem 4.1, Section missing 4).
Let ${\rm Mat}_{m,n}={\mathbb{C}}^{m\times n}$ be the space of $m\times n$ matrices with complex entries and consider the natural action of the torus $T=({\mathbb{C}}^{*})^{m}\times({\mathbb{C}}^{*})^{n}$ given by $(g,h)\cdot M=g\cdot M\cdot h^{-1}$ for all $M\in{\rm Mat}_{m,n}$ and $(g,h)\in T$. Let $X\subset{\rm Mat}_{m,n}$ be an irreducible $T$-subvariety and $C_{X}(t_{1},\ldots,t_{m},s_{1},\ldots,s_{n})$ be the polynomial representing the class $[X]^{T}$ of $X$ in $\text{H}_{T}^{\bullet}\left({\rm Mat}_{m,n}\right)={\mathbb{Z}}[t_{1},\ldots,t_{m},s_{1},\ldots,s_{n}]$. Then we have that $C_{X}(t_{1},\ldots,t_{m},-s_{1},\ldots,-s_{n})$ is a covolume polynomial.
We point out that the result of Theorem B is sharp in the sense that for arbitrary torus actions we can find simple instances where the equivariant class of an irreducible $T$-variety is represented by a polynomial which does not even have $M$-convex support (see Section missing 4).
We also connect the theory of Macaulay inverse systems over ${\mathbb{Z}}$ to the theory of Lorentzian polynomials. More precisely, we study the cohomology rings of smooth complex algebraic varieties over ${\mathbb{Z}}$ in terms of a suitable generic version of the Macaulay inverse system. In this characteristic-free situation over ${\mathbb{Z}}$, we show that the respective Macaulay dual generator is a denormalized Lorentzian polynomial (see Theorem C below).
Over a field of a characteristic zero, say the field of rational numbers $\mathbb{Q}$, the cohomology ring $\text{H}^{\bullet}(X,\mathbb{Q})$ of a smooth complex algebraic variety $X$ is a classical example of an Artinian graded Gorenstein algebra (due to the Poincaré duality of $\text{H}^{\bullet}(X,\mathbb{Q})$). Therefore, the Macaulay inverse system stands as a powerful tool to study the cohomology ring $\text{H}^{\bullet}(X,\mathbb{Q})$ because Artinian graded Gorenstein algebras can be defined in terms of the annihilator of a single polynomial in a dual polynomial ring (see, e.g., [Eis95, §21.2], [CLS11, §13.4]). Perhaps the most spectacular application of this idea is the result of Khovanskii and Pukhlikov [KP92] describing the cohomology ring of a smooth complete toric variety in terms of the annihilator of the volume polynomial. For more recent applications and related results, the reader is referred to [Kav11, HMM+13, KM21, KLM22, LMK22, MMSW22, HKM24].
Our interest is to study the cohomology ring $\text{H}^{\bullet}(X,{\mathbb{Z}})$ of a smooth complex algebraic variety $X$ as a ${\mathbb{Z}}$-algebra and to find suitable extensions of the aforementioned known results for $\text{H}^{\bullet}(X,\mathbb{Q})$. To substitute the fact that $\text{H}^{\bullet}(X,\mathbb{Q})$ is an Artinian graded Gorenstein algebra, we should have that the structure morphism $f\colon{\text{Spec}}\left(\text{H}^{\bullet}(X,{\mathbb{Z}})\right)\rightarrow{\text{Spec}}({\mathbb{Z}})$ is a finite Gorenstein morphism (i.e.; $f$ is a finite flat morphism and all its fibers are Gorenstein rings); if this happens, we say $\text{H}^{\bullet}(X,{\mathbb{Z}})$ is Artinian Gorenstein over ${\mathbb{Z}}$. We also need to consider generic versions of the Matlis and Macaulay dual functors. (These generic versions have become of interest recently, see, e.g., [Smi18, CCRS21, KK22, CR23].) Let $S={\mathbb{Z}}[x_{1},\ldots,x_{n}]$ be a positively graded polynomial ring and consider the inverse polynomial ring $T={\mathbb{Z}}[y_{1},\ldots,y_{n}]$ under the identification $y_{i}=x_{i}^{-1}$. For any homogeneous ideal $I\subset S$ such that the quotient $R=S/I$ is a finitely generated flat ${\mathbb{Z}}$-module, we have a successful Macaulay inverse system $I^{\perp_{{\mathbb{Z}}}}=\{G\in T\mid g\cdot G=0\text{ for all }g\in I\}$ that can be computed with the graded Matlis dual $R^{\vee_{\mathbb{Z}}}={}^{*}{\text{Hom}}_{\mathbb{Z}}(R,{\mathbb{Z}})$. We make the necessary developments of these notions in §missing 6.1. Our main result regarding cohomology rings is the following theorem.
Theorem C (Theorem 6.2, §missing 6.2).
Let $X$ be a $d$-dimensional smooth complex algebraic variety. Suppose that the cohomology ring $R=\bigoplus_{i=0}^{d}\text{H}^{2i}(X,{\mathbb{Z}})$ is a flat ${\mathbb{Z}}$-algebra (i.e., it is ${\mathbb{Z}}$-torsion-free). Let $\rho\colon R_{d}=\text{H}^{2d}(X,{\mathbb{Z}})\rightarrow{\mathbb{Z}}$ be the natural degree map. Choose a graded presentation $R\cong S/I$ where $S={\mathbb{Z}}[x_{1},\ldots,x_{n}]$, $\delta_{i}=\deg(x_{i})>0$, and $I\subset S$ is a homogeneous ideal. Let $\delta=\delta_{1}+\cdots+\delta_{n}$. Then the following statements hold:
  1. (i)
    $R$ is Artinian Gorenstein over ${\mathbb{Z}}$.
  2. (ii)
    We have the isomorphisms $\omega_{R/{\mathbb{Z}}}=\text{Ext}_{S}^{n}(R,S(-\delta))\cong{}^{*}{\text{Hom}}_{\mathbb{Z}}(R,{\mathbb{Z}})\cong R(d)$.
  3. (iii)
    Consider the inverse polynomial ring $T={\mathbb{Z}}[y_{1},\ldots,y_{n}]$ with the identification $y_{i}=x_{i}^{-1}$. Then the ideal $I\subset S$ is given as the annihilator
    $I\;=\;\{g\in S\mid g\cdot G_{R}=0\}$
    of the inverse polynomial
    $G_{R}(y_{1},\ldots,y_{n})\;=\;\sum_{\alpha_{1}\delta_{1}+\cdots+\alpha_{n}\delta_{n}=d}\rho\big{(}x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}\big{)}\,y_{1}^{\alpha_{1}}\cdots y_{n}^{\alpha_{n}}\;\;\in\;\;T={\mathbb{Z}}[y_{1},\ldots,y_{n}].$
  4. (iv)
    Assume also that $X$ is complete and that each $x_{i}$ is equal to the first Chern class $c_{1}(L_{i})$ of a nef line bundle $L_{i}$ on $X$. Then the normalization
    $N(G_{R})\;\in\;\mathbb{R}[y_{1},\ldots,y_{n}]$
    of $G_{R}$ is a Lorentzian polynomial.
  5. (v)
    Assume also that $X=X_{\Sigma}$ is a complete toric variety and that each $x_{i}$ is equal to the class of a torus-invariant nef Cartier divisor $D_{i}$ on $X$. Let $P_{i}=P_{D_{i}}$ be the polytope associated to $D_{i}$. Then
    $G_{R}(y_{1},\ldots,y_{n})\;=\;\sum_{\alpha_{1}+\cdots+\alpha_{n}=d}\mathrm{MV}_{\alpha}(P_{1},\ldots,P_{n})\,y_{1}^{\alpha_{1}}\cdots y_{n}^{\alpha_{n}}\;\;\in\;\;T={\mathbb{Z}}[y_{1},\ldots,y_{n}],$
    where $\mathrm{MV}_{\alpha}(P_{1},\ldots,P_{n})$ is the mixed volume of $P_{1},\ldots,P_{n}$ of type $\alpha=(\alpha_{1},\ldots,\alpha_{n})$.
Motivated by Theorem C, we say that $G_{R}\in T={\mathbb{Z}}[y_{1},\ldots,y_{n}]$ is the Macaulay dual generator of the cohomology ring $R=\text{H}^{\bullet}(X,{\mathbb{Z}})$ over ${\mathbb{Z}}$. We perform several explicit computations with examples in §missing 6.2.
Outline. The basic outline of this paper is as follows. In Section missing 2, we recall some basic results and we fix the notation. In Section missing 3, we study the multidegree polynomial of prime ideals in positive but not necessarily standard gradings. We prove Theorem B in Section missing 4. The proofs of Theorem A and Section missing 1 are given in Section missing 5. Finally, Section missing 6 contains the proof of Theorem C.
2. Preliminaries
In this section, we set up the notation used throughout the paper. We also present some preliminary results necessary to prove our main theorems.
2.1. Double Schubert polynomials
Let $p\geq 1$ be a positive integer. Let $[p]:=\{1,2,\dots,p\}$ and $\binom{[p]}{k}$ denote the set of size $k$ subsets of $[p]$. For $I=\{i_{1}<i_{2}<\cdots<i_{k}\}$, $J=\{j_{1}<j_{2}<\cdots<j_{k}\}\in\binom{[p]}{k}$, we define a partial order on $\binom{[p]}{k}$ where $I\leq J$ if $i_{m}\leq j_{m}$ for all $m\in[k]$. We write $S_{p}$ for the symmetric group of permutations on $[p]$. For any $w\in S_{p}$, we use one-line notation, i.e., $w=(a_{1},a_{2},\dots,a_{p})$ with $w(i)=a_{i}$. For any $k\in[p]$, we write $w[k]:=\{w(1),w(2),\dots,w(k)\}$. We say $(i,j)$ is an inversion of $w$ if $1\leq i<j\leq p$ and $w(i)>w(j)$. For every $i\in[p-1]=\{1,\ldots,p-1\}$, we have the transposition $\mathfrak{s}_{i}=(i,i+1)\in S_{p}$. The length $\ell(w)$ of a permutation $w$ is the number of inversions of $w$ or, equivalently, the minimum number of transpositions $\mathfrak{s}_{i}$ of which $w$ can be written as a product.
We equip $S_{p}$ with a partial order structure known as Bruhat order. For comprehensive references on Bruhat order and related topics, we point to [Hum90, BB05].
[…]
For $u,w\in S_{p}$, we say $w\geq u$ in Bruhat order if $w[i]\geq u[i]$ for all $i\in[p]$.
Double Schubert polynomials were introduced by Lascoux and Schützenberger [LS82] and can be defined via divided difference operators.
[…]
The $i$-th divided difference operator $\partial_{i}$ takes each polynomial $f\in{\mathbb{Z}}[t_{1},\dots,t_{p}]$ to
$\partial_{i}f(t_{1},\dots,t_{p})\;=\;\frac{f(t_{1},\dots,t_{p})-f(t_{1},\dots,t_{i-1},t_{i+1},t_{i},t_{i+2}\dots,t_{p})}{t_{i}-t_{i+1}}.$
The Schubert polynomial for $w\in S_{p}$ is defined by the recursion
$\mathfrak{S}_{w}(t_{1},\dots,t_{p})=\partial_{i}\mathfrak{S}_{w\mathfrak{s}_{i}}(t_{1},\dots,t_{p})$
whenever $\ell(w)<\ell(w\mathfrak{s}_{i})$, with initial data $\mathfrak{S}_{w_{0}}(t_{1},\ldots,t_{p})=\prod_{i=1}^{p-1}t_{i}^{p-i}\in{\mathbb{Z}}[t_{1},\dots,t_{p}]$. The double Schubert polynomials $\mathfrak{S}_{w}({\textbf{t}},{\textbf{s}})$ are defined by the same recursion where the divided difference operators only act on the $t_{i}$’s, but starting from $\mathfrak{S}_{w_{0}}({\textbf{t}},{\textbf{s}})=\prod_{i+j\leq p}(t_{i}-s_{j})\in{\mathbb{Z}}[t_{1},\dots,t_{p},s_{1},\dots,s_{p}]$.
2.2. Multidegrees
Here we recall the definition and basic properties of multidegrees as presented in [KM05] and [MS05, Chapter 8].
Let $R=\mathbb{k}[x_{1},\ldots,x_{n}]$ be a ${\mathbb{Z}}^{p}$-graded polynomial ring over a field $\mathbb{k}$. Let $M$ be a finitely generated ${\mathbb{Z}}^{p}$-graded $R$-module and $F_{\bullet}$ be a ${\mathbb{Z}}^{p}$-graded free $R$-resolution $F_{\bullet}\colon\;\cdots\rightarrow F_{i}\rightarrow F_{i-1}\rightarrow\cdots\rightarrow F_{1}\rightarrow F_{0}$ of $M$. Let $t_{1},\ldots,t_{p}$ be variables over ${\mathbb{Z}}$ and consider the Laurent polynomial ring ${\mathbb{Z}}[{\textbf{t}},\mathbf{t^{-1}}]={\mathbb{Z}}[t_{1},\ldots,t_{p},t_{1}^{-1},\ldots,t_{p}^{-1}]$, where the variable $t_{i}$ corresponds with the $i$-th elementary vector ${\mathbf{e}}_{i}\in{\mathbb{Z}}^{p}$. If we write $F_{i}=\bigoplus_{j}R(-\mathbf{b}_{i,j})$ with $\mathbf{b}_{i,j}=(\mathbf{b}_{i,j,1},\ldots,\mathbf{b}_{i,j,p})\in{\mathbb{Z}}^{p}$, then we define the Laurent polynomial $\left[F_{i}\right]_{\textbf{t}}\,:=\,\sum_{j}{\textbf{t}}^{\mathbf{b}_{i,j}}=\sum_{j}t_{1}^{\mathbf{b}_{i,j,1}}\cdots t_{p}^{\mathbf{b}_{i,j,p}}.$
[…]
The $K$-polynomial of $M$ is the Laurent polynomial given by
$\mathcal{K}(M;{\textbf{t}})\;:=\;\sum_{i}{(-1)}^{i}\left[F_{i}\right]_{\textbf{t}}.$
We have that, even if the grading of $R$ is non-positive and we do not have a well-defined notion of Hilbert series, the above definition of $K$-polynomial is an invariant of the module $M$ and it does not depend on the chosen free $R$-resolution $F_{\bullet}$ (see [MS05, Theorem 8.34]).
[…]
The multidegree polynomial of $M$ is the homogeneous polynomial $\mathcal{C}(M;{\textbf{t}})\in{\mathbb{Z}}[{\textbf{t}}]$ given as the sum of all terms in $\mathcal{K}(M;\mathbf{1}-{\textbf{t}})=\mathcal{K}(M;1-t_{1},\ldots,1-t_{p})$ having total degree ${\rm codim}(M)$, which is the lowest degree appearing.
2.3. Lorentzian and covolume polynomials
In this subsection, we briefly recall Lorentzian polynomials, dually Lorentzian polynomials, and covolume polynomials.
A subset $J\subset\mathbb{N}^{p}$ is called $\mathrm{M}$-convex if for any $q=(q_{1},\dots,q_{p})$ and $r=(r_{1},\dots,r_{p})$ in $J$, and any $i$ where $q_{i}<r_{i}$, there exists $j$ such that $q_{j}>r_{j}$ and both points $q+{\mathbf{e}}_{i}-{\mathbf{e}}_{j}$ and $r-{\mathbf{e}}_{i}+{\mathbf{e}}_{j}$ are also contained in $J$. We point out that M-convex sets are equivalent to basis elements of discrete polymatroids [Mur03, Sch03] and integer points of generalized permutohedra [Pos09]. A polynomial with M-convex support has saturated Newton polytope (SNP). In [MTY19], Monical, Tokcan, and Yong conjectured that many polynomials in algebraic combinatorics have saturated Newton polytopes.
Let $h(t_{1},\dots,t_{p})$ be a homogeneous polynomial of degree $d$ in $\mathbb{R}[{\textbf{t}}]=\mathbb{R}[t_{1},\ldots,t_{p}]$.
[…]
The homogeneous polynomial $h$ is called Lorentzian if the following conditions hold:
  1. (i)
    The coefficients of $h$ are nonnegative.
  2. (ii)
    The support of $h$ is M-convex.
  3. (iii)
    The quadratic form $\frac{\partial}{\partial t_{i_{1}}}\frac{\partial}{\partial t_{i_{2}}}\cdots\frac{\partial}{\partial t_{i_{e}}}h$ has at most one positive eigenvalue for any $i_{1},\ldots,i_{e}\in[p]$ where $e=d-2$.
There are several linear operators discussed in [BH20] that preserve the Lorentzian property. In particular, the normalization operator
$N\left(\sum_{{\mathbf{n}}}a_{\mathbf{n}}{\textbf{t}}^{\mathbf{n}}\right)\;:=\;\sum_{{\mathbf{n}}}\frac{a_{\mathbf{n}}}{{\mathbf{n}}!}{\textbf{t}}^{\mathbf{n}}\qquad\text{ where }\qquad{\mathbf{n}}!:=n_{1}!\cdots n_{p}!,$
preserves the Lorentzian property [BH20, Corollary 3.7].
In [Alu24], Aluffi defined the notion of covolume polynomials. These polynomials arise by considering the Chow classes of irreducible subvarieties of a product of projective spaces.
[…]
Let ${\mathbb{P}}={\mathbb{P}}_{\mathbb{k}}^{m_{1}}\times_{\mathbb{k}}\cdots\times_{\mathbb{k}}{\mathbb{P}}_{\mathbb{k}}^{m_{p}}$ be a multiprojective space over a field $\mathbb{k}$. Let $X\subset{\mathbb{P}}$ be an irreducible subvariety of codimension $c$. The class of $X$ can be written as
$[X]\;=\;\sum_{|{\mathbf{n}}|=c}a_{\mathbf{n}}\,H_{1}^{n_{1}}\cdots H_{p}^{n_{p}}\;\in\;A^{\bullet}({\mathbb{P}})={\mathbb{Z}}[H_{1},\dots,H_{p}]/\left(H_{1}^{m_{1}+1},\dots,H_{p}^{m_{p}+1}\right).$
Let $P_{[X]}(t_{1},\ldots,t_{p})=\sum_{|{\mathbf{n}}|=c}a_{\mathbf{n}}{\textbf{t}}^{{\mathbf{n}}}$ be the polynomial associated to the class $[X]$ of $X\subset{\mathbb{P}}$. A polynomial $P(t_{1},\dots,t_{p})\in\mathbb{R}[t_{1},\ldots,t_{p}]$ with nonnegative real coefficients is a covolume polynomial if it is a limit of polynomials of the form $cP_{[X]}$ for a positive real number $c$ and an irreducible subvariety $X$ of ${\mathbb{P}}$.
We are interested in the family of dually Lorentzian polynomials introduced by Ross, Süß, and Wannerer [RSW23].
[…]
A polynomial $h\in\mathbb{R}[t_{1},\dots,t_{p}]$ is dually Lorentzian if
$N\left(t_{1}^{m_{1}}\cdots t_{p}^{m_{p}}h\left(t_{1}^{-1},\dots,t_{p}^{-1}\right)\right)$
is Lorentzian for some ${\mathbf{m}}=(m_{1},\ldots,m_{p})$ sufficiently large (i.e., such that $t_{1}^{m_{1}}\cdots t_{p}^{m_{p}}h\left(t_{1}^{-1},\dots,t_{p}^{-1}\right)$ is a polynomial).
As shown by Aluffi [Alu24, Proposition 2.8], covolume polynomials form a subfamily of the family of dually Lorentzian polynomials.
[…]
For any monomial $x_{1}^{\alpha_{1}}\cdots x_{p}^{\alpha_{n}}$ and any polynomial $h\in\mathbb{R}[t_{1},\ldots,t_{p}]$, we have that $h$ is dually Lorentzian if and only if $x_{1}^{\alpha_{1}}\cdots x_{p}^{\alpha_{n}}h$ is dually Lorentzian.
Proof.
This follows directly from [HMMSD22, Lemma 7]. ∎
Finally, we introduce the following definition.
[…]
For any polynomial $h=\sum_{\mathbf{n}}a_{\mathbf{n}}{\textbf{t}}^{\mathbf{n}}\in\mathbb{R}[t_{1},\ldots,t_{p}]$ and any ${\mathbf{w}}\in\mathbb{N}^{p}$, the ${\mathbf{w}}$-truncation of $h$ is the polynomial given by $\sum_{{\mathbf{n}}\leq{\mathbf{w}}}a_{\mathbf{n}}{\textbf{t}}^{\mathbf{n}}\in\mathbb{R}[t_{1},\ldots,t_{p}].$
3. Multidegree polynomials of prime ideals in non-standard gradings
In this section, working over an arbitrary positive $\mathbb{N}^{p}$-grading, we show that the multidegree polynomial of a prime ideal is a covolume polynomial. Our main tool is the technique of standardization that was used in [CCRMMn23, CCRC23].
Let $\mathbb{k}$ be an algebraically closed field and $R=\mathbb{k}[x_{1},\ldots,x_{n}]$ be a positively $\mathbb{N}^{p}$-graded polynomial ring (that is, $\deg(x_{i})\in\mathbb{N}^{p}\setminus\{\mathbf{0}\}$ for all $1\leq i\leq n$ and $\deg(\alpha)=\mathbf{0}\in\mathbb{N}^{p}$ for all $\alpha\in\mathbb{k}$). Let $M$ be a finitely generated ${\mathbb{Z}}^{p}$-graded $R$-module. Since $R$ is assumed to be positively graded, there is a well-defined notion of Hilbert series
$\text{Hilb}_{M}({\textbf{t}})\;=\;\text{Hilb}_{M}(t_{1},\ldots,t_{p})\;:=\;\sum_{{\mathbf{n}}\in{\mathbb{Z}}^{p}}\dim_{\mathbb{k}}\left(\left[M\right]_{\mathbf{n}}\right){\textbf{t}}^{\mathbf{n}}=\sum_{{\mathbf{n}}=(n_{1},\ldots,n_{p})\in{\mathbb{Z}}^{p}}\dim_{\mathbb{k}}\left(\left[M\right]_{\mathbf{n}}\right)t_{1}^{n_{1}}\cdots t_{p}^{n_{p}},$
and then we can write
$\text{Hilb}_{M}({\textbf{t}})\;=\;\frac{\mathcal{K}(M;{\textbf{t}})}{\prod_{i=1}^{n}\left(1-{\textbf{t}}^{\deg(x_{i})}\right)}.$
Below we discuss the case of a standard multigrading. This case is of special importance since it deals with closed subschemes of a product of projective spaces.
[…]
Assume that $R$ has a standard $\mathbb{N}^{p}$-grading (i.e., $|\deg(x_{i})|=1$) and that $R$ is the coordinate ring of ${\mathbb{P}}={\mathbb{P}}_{\mathbb{k}}^{m_{1}}\times_{\mathbb{k}}\cdots\times_{\mathbb{k}}{\mathbb{P}}_{\mathbb{k}}^{m_{p}}$. Let $X\subset{\mathbb{P}}$ be a $d$-dimensional integral closed subscheme with coordinate ring $R/P$, where $P$ is an $R$-homogeneous prime ideal. The class of $X$ in the Chow ring of ${\mathbb{P}}$ is given by
$\left[X\right]\;=\;\sum_{\begin{subarray}{c}{\mathbf{n}}=(n_{1},\ldots,n_{p})\in\mathbb{N}^{p}\\ |{\mathbf{n}}|=n_{1}+\cdots+n_{p}=d\end{subarray}}\deg_{\mathbb{P}}^{\mathbf{n}}(X)\cdot H_{1}^{m_{1}-n_{1}}\cdots H_{p}^{m_{p}-n_{p}}\;\in\;A^{\bullet}({\mathbb{P}})=\frac{{\mathbb{Z}}[H_{1},\ldots,H_{p}]}{\big{(}H_{1}^{m_{1}+1},\ldots,H_{p}^{m_{p}+1}\big{)}},$
where $H_{i}$ represents the class of the inverse image of a hyperplane of ${\mathbb{P}}_{\mathbb{k}}^{m_{i}}$. We say that $\deg_{\mathbb{P}}^{\mathbf{n}}(X)$ is the multidegree of $X$ of type ${\mathbf{n}}$. Then the multidegree polynomial of $R/P$ is given by
$\mathcal{C}(R/P;{\textbf{t}})\;=\;\sum_{{\mathbf{n}}\in\mathbb{N}^{p},\;|{\mathbf{n}}|=d}\deg_{\mathbb{P}}^{\mathbf{n}}(X)\cdot t_{1}^{m_{1}-n_{1}}\cdots t_{p}^{m_{p}-n_{p}}.$
The volume polynomial of $X$ (see [BH20, §4.2]) is given by
${\rm vol}_{X}({\textbf{t}})\;=\;\int\big{(}H_{1}t_{1}+\cdots+H_{p}t_{p}\big{)}^{d}\cap\left[X\right]\;=\;\sum_{{\mathbf{n}}\in\mathbb{N}^{p},\;|{\mathbf{n}}|=d}\deg_{\mathbb{P}}^{\mathbf{n}}(X)\cdot\frac{d!}{n_{1}!\cdots n_{p}!}\cdot t_{1}^{n_{1}}\cdots t_{p}^{n_{p}}.$
Due to a fundamental result of Brändén and Huh (see [BH20, Theorem 4.6]), we have that ${\rm vol}_{X}({\textbf{t}})$ is a Lorentzian polynomial.
The following lemma tells us that multidegree polynomials in a standard multigrading are dually Lorentzian [RSW23]. This result already appeared in [Alu24, Proposition 2.8] (also, see [HMMSD22, Theorem 6]), but here we give a short self-contained proof.
[…]
Keep the same notations and assumptions of Section missing 3. Consider the polynomial
$F(t_{1},\ldots,t_{p})\;=\;t_{1}^{m_{1}}\cdots t_{p}^{m_{p}}\cdot\mathcal{C}\Big{(}R/P;\frac{1}{t_{1}},\ldots,\frac{1}{t_{p}}\Big{)}.$
Then the normalization $N(F)$ is a Lorentzian polynomial (i.e., $\mathcal{C}(R/P;{\textbf{t}})$ is dually Lorentzian).
Proof.
From Section missing 3, we derive the following equalities
$N(F({\textbf{t}}))$ $\;=\;N\Bigg{(}t_{1}^{m_{1}}\cdots t_{p}^{m_{p}}\cdot\mathcal{C}\Big{(}R/P;\frac{1}{t_{1}},\ldots,\frac{1}{t_{p}}\Big{)}\Bigg{)}$
$\;=\;N\Bigg{(}t_{1}^{m_{1}}\cdots t_{p}^{m_{p}}\cdot\sum_{{\mathbf{n}}\in\mathbb{N}^{p},\;|{\mathbf{n}}|=d}\deg_{\mathbb{P}}^{\mathbf{n}}(X)\cdot\frac{1}{t_{1}^{m_{1}-n_{1}}}\cdots\frac{1}{t_{p}^{m_{p}-n_{p}}}\Bigg{)}$
$\;=\;N\Bigg{(}\sum_{{\mathbf{n}}\in\mathbb{N}^{p},\;|{\mathbf{n}}|=d}\deg_{\mathbb{P}}^{\mathbf{n}}(X)\cdot{t_{1}^{n_{1}}}\cdots{t_{p}^{n_{p}}}\Bigg{)}$
$\;=\;\sum_{{\mathbf{n}}\in\mathbb{N}^{p},\;|{\mathbf{n}}|=d}\deg_{\mathbb{P}}^{\mathbf{n}}(X)\cdot\frac{1}{n_{1}!\cdots n_{p}!}\cdot t_{1}^{n_{1}}\cdots t_{p}^{n_{p}}$
$\;=\;\frac{1}{d!}\cdot{\rm vol}_{X}({\textbf{t}}).$
Therefore [BH20, Theorem 4.6] implies that $N(F)$ is Lorentzian. ∎
We now describe some basic properties of the process of standardization as developed in [CCRMMn23, CCRC23]. For the rest of the section, the following setup is fixed.
[…]
For $1\leq i\leq n$, let $\ell_{i}=|\deg(x_{i})|$ be the total degree of the variable $x_{i}$. Let
$S\;=\;\mathbb{k}\left[y_{i,j}\mid 1\leq i\leq n\text{ and }1\leq j\leq\ell_{i}\right]$
be a standard $\mathbb{N}^{p}$-graded polynomial ring such that
$\deg(x_{i})\;=\;\sum_{j=1}^{\ell_{i}}\deg(y_{i,j})\quad\text{for all \;\;$1\leq i\leq n$.}$
We define the $\mathbb{N}^{p}$-graded $\mathbb{k}$-algebra homomorphism
$\phi\colon R=\mathbb{k}[\mathbf{x}]\longrightarrow S=\mathbb{k}[\mathbf{y}],\quad\phi(x_{i})=y_{i,1}y_{i,2}\cdots y_{i,\ell_{i}}.$
For an $R$-homogeneous ideal $I\subset R$, we say that the extension $J=\phi(I)S$ is the standardization of $I$, since $J\subset S$ is an $S$-homogeneous ideal in the standard $\mathbb{N}^{p}$-graded polynomial ring $S$. By a slight abuse of notation, we consider both multidegree polynomials $\mathcal{C}(R/I;{\textbf{t}})$ and $\mathcal{C}(S/J;{\textbf{t}})$ as elements of the same polynomial ring ${\mathbb{Z}}[{\textbf{t}}]={\mathbb{Z}}[t_{1},\ldots,t_{p}]$.
The next theorem contains some of the basic and desirable properties of the standardization of an ideal.
Theorem 3.1 ([CCRC23, Theorem 7.2], [CCRMMn23, Proposition 4.2]).
Let $I\subset R$ be an $R$-homogeneous ideal and $J=\phi(I)S$ be its standardization. Then the following statements hold:
  1. (i)
    ${\rm codim}(I)={\rm codim}(J)$.
  2. (ii)
    $I\subset R$ and $J\subset S$ have the same $\mathbb{N}^{p}$-graded Betti numbers.
  3. (iii)
    $\mathcal{K}(R/I;{\textbf{t}})=\mathcal{K}(S/J;{\textbf{t}})$ and $\mathcal{C}(R/I;{\textbf{t}})=\mathcal{C}(S/J;{\textbf{t}})$.
  4. (iv)
    $R/I$ is a Cohen–Macaulay ring if and only if $S/J$ is a Cohen–Macaulay ring.
  5. (v)
    Let $>$ be a monomial order on $R$ and $>^{\prime}$ be a monomial order on $S$ which is compatible with $\phi$ (i.e., if $f,g\in R$ with $f>g$, then $\phi(f)>^{\prime}\phi(g)$). Then ${\rm in}_{>^{\prime}}(J)=\phi({\rm in}_{>}(I))S$.
  6. (vi)
    If $I\subset R$ is a prime ideal and it does not contain any variable, then $J\subset S$ is also a prime ideal.
Finally, we are ready to present the main result of this section. It yields a large new family of covolume polynomials.
Theorem 3.2.
Let $P\subset R$ be a prime $R$-homogeneous ideal. Then $\mathcal{C}(R/P;{\textbf{t}})$ is a covolume polynomial.
Proof.
Let =$\mathcal{L}=\{i\mid x_{i}\in P\}$ be the set that indexes the variables belonging to $P$. Write $P=P^{\prime}+\left(x_{i}\mid i\in\mathcal{L}\right)$ with $P^{\prime}\subset R$ only involving the variables not in $\mathcal{L}$. Let $T=R\left[z_{i}\mid i\in\mathcal{L}\right]$ be a positively $\mathbb{N}^{p}$-graded polynomial ring extending the grading of $R$ and with $\deg(z_{i})=\deg(x_{i})$. Let $\mathfrak{P}=P^{\prime}T+\left(x_{i}-z_{i}\mid i\in\mathcal{L}\right)\subset T$. Notice that $\mathfrak{P}$ is a prime ideal containing no variable and that
$\mathcal{C}(T/\mathfrak{P};{\textbf{t}})\;=\;\prod_{i\in\mathcal{L}}\langle\deg(x_{i}-z_{i}),{\textbf{t}}\rangle\cdot\mathcal{C}(T/P^{\prime}T;{\textbf{t}})\;=\;\prod_{i\in\mathcal{L}}\langle\deg(x_{i}),{\textbf{t}}\rangle\cdot\mathcal{C}(R/P^{\prime};{\textbf{t}})\;=\;\mathcal{C}(R/P;{\textbf{t}}),$
where $\langle\deg(x_{i}),{\textbf{t}}\rangle=a_{i,1}t_{1}+\cdots+a_{i,p}t_{p}\in\mathbb{N}[{\textbf{t}}]$ after writing $(a_{i,1},\ldots,a_{i,p})=\deg(x_{i})\in\mathbb{N}^{p}$ (see [MS05, Exercise 8.12]). Let $W=S\left[w_{i,j}\mid i\in\mathcal{L},1\leq j\leq\ell_{i}\right]$ be a standard $\mathbb{N}^{p}$-graded polynomial ring where we consider the corresponding standardization $\mathfrak{Q}\subset W$ of $\mathfrak{P}\subset T$. From Theorem 3.1(iii), we get
$\mathcal{C}(R/P;{\textbf{t}})\;=\;\mathcal{C}(T/\mathfrak{P};{\textbf{t}})\;=\;\mathcal{C}(W/\mathfrak{Q};{\textbf{t}}).$
Finally, $\mathfrak{Q}\subset W$ is a prime ideal by Theorem 3.1(vi), and so the result follows by Section missing 3. ∎
4. Equivariant cohomology in multigraded commutative algebra
In this section, we study the equivariant cohomology of the irreducible varieties that appear in multigraded commutative algebra. Here we show that, under suitable conditions, equivariant classes of irreducible multigraded varieties yield covolume polynomials (up to changing the sign of negative coefficients). We follow the references [AF24] and [CG10, Chapter 5] for the basics of equivariant cohomology and equivariant $K$-theory.
Let $R={\mathbb{C}}[x_{1},\ldots,x_{n}]$ be a polynomial ring with a ${\mathbb{Z}}^{p}$-grading. Let $T=({\mathbb{C}}^{*})^{p}$ be a torus acting on $\mathbb{A}_{\mathbb{C}}^{n}$. Any such ${\mathbb{Z}}^{p}$-grading can be specified by the torus weights. More precisely, if we have the action
$\left(h_{1},\ldots,h_{p}\right)\cdot\left(a_{1},\ldots,a_{n}\right)\;=\;\big{(}h_{1}^{d_{1,1}}h_{2}^{d_{1,2}}\cdots h_{p}^{d_{1,p}}a_{1},\,h_{1}^{d_{2,1}}h_{2}^{d_{2,2}}\cdots h_{p}^{d_{2,p}}a_{2},\,\ldots,\,h_{1}^{d_{n,1}}h_{2}^{d_{n,2}}\cdots h_{p}^{d_{n,p}}a_{n}\big{)}$
for all $(h_{1},\ldots,h_{p})\in T$ and $(a_{1},\ldots,a_{n})\in\mathbb{A}_{\mathbb{C}}^{n}$, then we obtain the specific grading $\deg(x_{i})=\mathbf{d}_{i}=(d_{i,1},d_{i,2},\ldots,d_{i,p})\in{\mathbb{Z}}^{p}$ for all $1\leq i\leq n$. There is a bijective correspondence between ${\mathbb{Z}}^{p}$-graded $R$-modules and $T$-equivariant coherent sheaves on $\mathbb{A}_{\mathbb{C}}^{n}$. Since $\mathbb{A}_{\mathbb{C}}^{n}$ is smooth, the Grothendieck ring of $T$-equivariant vector bundles coincides with the Grothendieck ring of $T$-equivariant coherent sheaves. We denote this ring by $K_{T}(\mathbb{A}_{\mathbb{C}}^{n})$. The Grothendieck ring $K_{T}(\mathbb{A}_{\mathbb{C}}^{n})$ coincides with the Grothendieck ring of a point and the representation ring of $T$:
$K_{T}(\mathbb{A}_{\mathbb{C}}^{n})\;\cong\;K_{T}({\rm pt})\;\cong\;R(T)\;=\;{\mathbb{Z}}[t_{1},\ldots,t_{p},t_{1}^{-1},\ldots,t_{p}^{-1}].$
Given a finitely generated ${\mathbb{Z}}^{p}$-graded $R$-module $M$ (i.e., $\widetilde{M}$ is a $T$-equivariant coherent sheaf on $\mathbb{A}_{\mathbb{C}}^{n}$), we denote by $[M]^{T}$ the equivariant class of $M$ in $K_{T}(\mathbb{A}_{\mathbb{C}}^{n})\cong{\mathbb{Z}}[t_{1},\ldots,t_{p},t_{1}^{-1},\ldots,t_{p}^{-1}]$.
We consider the $T$-equivariant cohomology ring
$\text{H}_{T}^{\bullet}(\mathbb{A}_{\mathbb{C}}^{n})\;:=\;\text{H}^{\bullet}\big{(}\mathbb{E}T\times^{T}\mathbb{A}_{\mathbb{C}}^{n}\big{)},$
where $\mathbb{E}T$ is contractible with $T$ acting freely (on the right). Then $\mathbb{B}T:=\mathbb{E}T/T$ is a classifying space for $T$. Since we can take $\mathbb{E}T=\left({\mathbb{C}}^{\infty}\setminus\{0\}\right)^{n}$ and $\mathbb{B}T=\left({\mathbb{P}}_{\mathbb{C}}^{\infty}\right)^{n}$, it follows that
$\text{H}_{T}^{\bullet}(\mathbb{A}_{\mathbb{C}}^{n})\;\cong\;\Lambda_{T}:=\text{H}^{\bullet}_{T}({\rm pt})\;=\;\text{H}^{\bullet}(\mathbb{B}T)\;=\;{\mathbb{Z}}[t_{1},\ldots,t_{p}].$
Given a $T$-subvariety $X\subset\mathbb{A}_{\mathbb{C}}^{n}$, we denote by $[X]^{T}:=[\mathbb{E}T\times^{T}X]$ the equivariant class of $X$ in $\text{H}_{T}^{\bullet}(\mathbb{A}_{\mathbb{C}}^{n})\cong{\mathbb{Z}}[t_{1},\ldots,t_{p}]$.
The following remark allows us to apply our results in Section missing 3 to equivariant cohomology.
[…]
Let $M$ be a finitely generated ${\mathbb{Z}}^{p}$-graded $R$-module and $X\subset\mathbb{A}_{\mathbb{C}}^{n}$ be a $T$-subvariety with coordinate ring $R/I$. Let $K_{M}(t_{1},\ldots,t_{p})$ and $C_{X}(t_{1},\ldots,t_{p})$ be the Laurent polynomial and the polynomial representing the classes $[M]^{T}\in K_{T}(\mathbb{A}_{\mathbb{C}}^{n})$ and $[X]^{T}\in\text{H}_{T}^{\bullet}(\mathbb{A}_{\mathbb{C}}^{n})$, respectively. Then we have the equalities
$K_{M}(t_{1},\ldots,t_{p})\;=\;\mathcal{K}(M;t_{1},\ldots,t_{p})\qquad\text{ and }\qquad C_{X}(t_{1},\ldots,t_{p})\;=\;\mathcal{C}(R/I;t_{1},\ldots,t_{p}).$
Proof.
See, e.g., [KM05, §] or [KMS06, Proposition 1.19]. ∎
Then Theorem 3.2 poses the question of whether the class $[X]^{T}$ of any irreducible $T$-subvariety $X\subset\mathbb{A}_{\mathbb{C}}^{n}$ yields a covolume polynomial. The following example shows that it is not always the case for arbitrary $T$-actions.
[…]
Consider $\mathbb{A}_{\mathbb{C}}^{2}$ and the torus $T=({\mathbb{C}}^{*})^{2}$ with the action
$(h_{1},h_{2})\cdot(a_{1},a_{2})\;:=\;(h_{1}h_{2}a_{1},h_{1}h_{2}^{-1}a_{2})$
for all $(a_{1},a_{2})\in\mathbb{A}_{\mathbb{C}}^{2}$ and $(h_{1},h_{2})\in T$. Let $R={\mathbb{C}}[x,y]$ be the coordinate ring of $\mathbb{A}_{\mathbb{C}}^{2}$. The $T$-action above specifies the grading $\deg(x)=(1,1)\in{\mathbb{Z}}^{2}$ and $\deg(y)=(1,-1)\in{\mathbb{Z}}^{2}$. Consider the variety $X=\{(0,0)\}\subset\mathbb{A}_{\mathbb{C}}^{2}$ given by the origin. It is clearly irreducible, but the polynomial representative of the cohomology class
[…]
is not a covolume polynomial (its support is not $M$-convex) after changing signs.
For the rest of this section, we shall use the following setup that avoids the complications of the previous example.
[…]
Assume that the torus $T$ is given as $T=({\mathbb{C}}^{*})^{q}\times({\mathbb{C}}^{*})^{p-q}$ where “$({\mathbb{C}}^{*})^{q}$ comes with positive weights and $({\mathbb{C}}^{*})^{p-q}$ comes with negative weights”. More precisely, we require that
$\mathbf{d}_{i}\in\mathbb{N}^{p}\setminus\{\mathbf{0}\}\quad\text{ for all $1\leq i\leq q$}$
and
$-\mathbf{d}_{i}\in\mathbb{N}^{p}\setminus\{\mathbf{0}\}\quad\text{ for all $q+1\leq i\leq p$}.$
In this case we say that the action of $T$ determines a twisted positive grading on $R$.
The following lemma tells us that we can “flip” twisted positive gradings to positive gradings. A version of this lemma appeared in [CCRMMn23, Lemma 3.3]. Let $\widetilde{R}=\mathbb{k}[x_{1},\ldots,x_{n}]$ be a polynomial ring with the same variables as $R$ but with grading given by setting $\deg(x_{i})=\mathbf{d}_{i}$ for $1\leq i\leq q$ and $\deg(x_{i})=-\mathbf{d}_{i}$ for $q+1\leq i\leq p$. Notice that $\widetilde{R}$ has a positive $\mathbb{N}^{p}$-grading.
[…]
Assume Section missing 4. Let $I\subset R$ be a ${\mathbb{Z}}^{p}$-graded ideal, and also denote by $I$ the corresponding $\mathbb{N}^{p}$-graded ideal in $\widetilde{R}$. Then we have $\mathcal{C}(\widetilde{R}/I;t_{1},\ldots,t_{q},t_{q+1},\ldots,t_{p})=\mathcal{C}(R/I;t_{1},\ldots,t_{q},-t_{q+1},\ldots,-t_{p}).$
Proof.
Let $r=p-q$ and set $s_{i}=t_{p+i}$ for $1\leq i\leq r$. Notice that, if $\widetilde{F}_{\bullet}$ is a $({\mathbb{Z}}^{q}\oplus{\mathbb{Z}}^{r})$-graded free $\widetilde{R}$-resolution of $\widetilde{R}/I$ with $\widetilde{F}_{i}=\bigoplus_{j}\widetilde{R}(-\mathbf{a}_{i,j},-\mathbf{b}_{i,j})$, then there is a corresponding $({\mathbb{Z}}^{q}\oplus{\mathbb{Z}}^{r})$-graded free $R$-resolution $F_{\bullet}$ of $R/I$ with $F_{i}=\bigoplus_{j}R(-\mathbf{a}_{i,j},\mathbf{b}_{i,j})$. By definition, this yields the equality of $K$-polynomials
$\mathcal{K}(\widetilde{R}/I;{\textbf{t}},{\textbf{s}})=\mathcal{K}(\widetilde{R}/I;t_{1},\ldots,t_{q},s_{1},\ldots,s_{r})=\mathcal{K}(R/I;t_{1},\ldots,t_{q},s_{1}^{-1},\ldots,s_{r}^{-1})=\mathcal{K}(R/I;{\textbf{t}},{\textbf{s}}^{\mathbf{-1}}).$
From [MS05, Claim 8.54], we have $\mathcal{K}(R/I;\mathbf{1-t},\mathbf{1-s})=\mathcal{C}(R/I;{\textbf{t}},{\textbf{s}})+Q({\textbf{t}},{\textbf{s}})$, where $Q({\textbf{t}},{\textbf{s}})$ is the sum of the terms of degree at least ${\rm codim}(I)+1$. Equivalently, we get $\mathcal{K}(R/I;{\textbf{t}},{\textbf{s}})=\mathcal{C}(R/I;\mathbf{1-t},\mathbf{1-s})+Q(\mathbf{1-t},\mathbf{1-s})$. It then follows that
$\mathcal{K}(\widetilde{R}/I;\mathbf{1-t},\mathbf{1-s})=\mathcal{C}(R/I;t_{1},\ldots,t_{q},1-\tfrac{1}{1-s_{1}},\ldots,1-\tfrac{1}{1-s_{r}})+Q(t_{1},\ldots,t_{q},1-\tfrac{1}{1-s_{1}},\ldots,1-\tfrac{1}{1-s_{r}}).$
By expanding the right hand side of the above equality, the result of the lemma is obtained. ∎
We are now ready to present the main result of this section.
Theorem 4.1.
Assume Section missing 4. Let $X\subset\mathbb{A}_{\mathbb{C}}^{n}$ be an irreducible $T$-subvariety and $C_{X}(t_{1},\ldots,t_{p})$ be the polynomial representing the class $[X]^{T}\in\text{H}_{T}^{\bullet}(\mathbb{A}_{\mathbb{C}}^{n})$. Then $C_{X}(t_{1},\ldots,t_{q},-t_{q+1},\ldots,-t_{p})$ is a covolume polynomial.
Proof.
The result follows by combining Section missing 4, Section missing 4, and Theorem 3.2. ∎
Our main application of the above theorem is the following corollary.
[…]
Let ${\rm Mat}_{m,n}={\mathbb{C}}^{m\times n}$ be the space of $m\times n$ matrices with complex entries and consider the natural action of the torus $T=({\mathbb{C}}^{*})^{m}\times({\mathbb{C}}^{*})^{n}$ given by $(g,h)\cdot M=g\cdot M\cdot h^{-1}$ for all $M\in{\rm Mat}_{m,n}$ and $(g,h)\in T$. Let $X\subset{\rm Mat}_{m,n}$ be an irreducible $T$-variety and $C_{X}(t_{1},\ldots,t_{m},s_{1},\ldots,s_{n})$ be the polynomial representing the class $[X]^{T}$ in $\text{H}_{T}^{\bullet}\left({\rm Mat}_{m,n}\right)={\mathbb{Z}}[t_{1},\ldots,t_{m},s_{1},\ldots,s_{n}]$. Then we have that $C_{X}(t_{1},\ldots,t_{m},-s_{1},\ldots,-s_{n})$ is a covolume polynomial.
5. Equivariant cohomology of matrix Richardson varieties
In this section, we study the equivariant cohomology of matrix Richardson varieties for a pair of permutations. An interesting outcome of our approach is the definition of a new family of polynomials that we call double Richardson polynomials. These polynomials specialize to many polynomials of interest.
We take the viewpoint of studying Schubert varieties and Schubert polynomials through a certain degeneracy loci problem that was introduced by Fulton [Ful92]. Let $E_{\bullet}$ and $F_{\bullet}$ be two flagged complex vector spaces of dimension $n$ with
$F_{1}\subset\cdots\subset F_{n}=F\qquad\text{ and }\qquad E=E_{n}\twoheadrightarrow\cdots\twoheadrightarrow E_{1},$
where $\dim(F_{i})=\dim(E_{i})=i$. By choosing bases, we identify ${\text{Hom}}(F,E)$ with the space ${\rm Mat}_{n,n}\cong{\mathbb{C}}^{n\times n}$ of $n\times n$ matrices with complex entries. Consider the natural action of the torus $T=({\mathbb{C}}^{*})^{n}\times({\mathbb{C}}^{*})^{n}$ given by $(g,h)\cdot M=g\cdot M\cdot h^{-1}$ for all $M\in{\rm Mat}_{n,n}$ and $(g,h)\in T$. For a given permutation $w\in S_{n}$, we consider the degeneracy locus
$D_{w}\;:=\;\big{\{}\varphi\in{\text{Hom}}(F,E)\mid\text{rank }(\varphi_{p,q})\leq\text{rank }(w_{p,q})\text{ for all $1\leq p,q\leq n$}\big{\}},$
[…]
Choose a basis $\{f_{1},\ldots,f_{n}\}$ of $F$ such that $\{f_{1},\ldots,f_{i}\}$ is a basis of $F_{i}$ for all $1\leq i\leq n$. Let $\widetilde{F}_{\bullet}$ be the opposite flag
$\widetilde{F}_{1}\subset\cdots\subset\widetilde{F}_{n}\qquad\text{ where }\qquad\widetilde{F}_{i}=\langle f_{n-i+1},\ldots,f_{n}\rangle.$
For a given permutation $w\in S_{n}$, we consider the degeneracy locus
$D^{w}\;:=\;\big{\{}\varphi\in{\text{Hom}}(F,E)\mid\text{rank }(\varphi^{p,q})\leq\text{rank }(w^{p,q})\text{ for all $1\leq p,q\leq n$}\big{\}},$
[…]
A result of fundamental importance for us is the following geometric interpretation of double Schubert polynomials.
Theorem 5.1 ([FR03, KM05, AF24]).
The equivariant class $\left[D_{w}\right]^{T}$ of the matrix Schubert variety $D_{w}$ in the cohomology ring $\text{H}_{T}^{\bullet}({\rm Mat}_{n,n})\cong{\mathbb{Z}}[t_{1},\ldots,t_{n},s_{1},\ldots,s_{n}]$ is given by the double Schubert polynomial $\mathfrak{S}_{w}({\textbf{t}},{\textbf{s}})$.
The next lemma expresses the equivariant class of the degeneracy locus $D^{w}$.
[…]
The equivariant class $\left[D^{w}\right]^{T}\in\text{H}_{T}^{\bullet}({\rm Mat}_{n,n})$ of $D^{w}$ is given by the double Schubert polynomial
$\mathfrak{S}_{w_{0}w}(t_{1},\ldots,t_{n},s_{n},\ldots,s_{1})$
of $w_{0}w$ with a reverse ordering of the variables $s_{1},\ldots,s_{n}$.
Proof.
Let $\alpha\colon F\rightarrow F$ be the involution given by setting $\alpha(f_{i})=f_{n+1-i}$. To simplify notation, let $\mathcal{M}={\text{Hom}}(F,E)$. Let $\eta\colon T\rightarrow T$ be the involution given by $\eta(g_{1},\ldots,g_{n},h_{1},\ldots,h_{n})=(g_{1},\ldots,g_{n},h_{n},\ldots,h_{1})$. Then the induced involution $f\colon\mathcal{M}\rightarrow\mathcal{M},\varphi\mapsto\varphi\circ\alpha$ is an equivariant map with respect to $\eta$ because
$f\left((g,h)\cdot\varphi\right)\;=\;\eta(g,h)\cdot f(\varphi)$
for all $\varphi\in\mathcal{M}$ and $(g,h)\in T$. Then the induced pullback homomorphism is given by
$f^{*}\colon\;\text{H}_{T}^{\bullet}(\mathcal{M})={\mathbb{Z}}[{\textbf{t}},{\textbf{s}}]\;\rightarrow\;\text{H}_{T}^{\bullet}(\mathcal{M})={\mathbb{Z}}[{\textbf{t}},{\textbf{s}}],\qquad t_{i}\mapsto t_{i}\text{\; and \;}s_{j}\mapsto s_{n+1-j}$
(see [AF24, Exercise 3.2.3]). Let $\mathbb{E}_{m}=({\mathbb{C}}^{m}\setminus 0)^{n}\times({\mathbb{C}}^{m}\setminus 0)^{n}$. Then $\text{H}_{T}^{i}(\mathcal{M})=\text{H}^{i}(\mathbb{E}_{m}\times^{T}\mathcal{M})$ for $i<2m-1$ and the pullback map $f^{*}$ can be constructed by taking the pullback in (usual) cohomology along the equivariant map
$\mathbb{E}_{m}\times^{T}\mathcal{M}\;\xrightarrow{(\gamma,f)}\;\mathbb{E}_{m}\times^{T}\mathcal{M}$
where $\gamma\colon\mathbb{E}_{m}\rightarrow\mathbb{E}_{m}$ is the involution given by $\gamma(u_{1},\ldots,u_{n},v_{1},\ldots,v_{n})=(u_{1},\ldots,u_{n},v_{n},\ldots,v_{1})$ for all $(u_{1},\ldots,u_{n},v_{1},\ldots,v_{n})\in\mathbb{E}_{m}$ (see [AF24, Exercise 3.2.1]). Since $\mathbb{E}_{m}\times^{T}\mathcal{M}$ is smooth, it follows that
$f^{*}\left(\big{[}D_{w_{0}w}\big{]}^{T}\right)\;=\;\big{[}f^{-1}(D_{w_{0}w})\big{]}^{T}=\big{[}D^{w}\big{]}^{T}$
(see, e.g., [AF24, Proposition A.3.2], [Ful97, §B.3]). The result of the lemma now follows by invoking Theorem 5.1. ∎
We are now ready to introduce the two main objects of this section.
[…]
For a pair of permutations $(w,u)$ in $S_{n}$, we have:
  1. (i)
    $D_{u}^{w}:=D_{u}\cap D^{w}$ is the matrix Richardson variety.
  2. (ii)
    $\mathfrak{R}_{w/u}({\textbf{t}},{\textbf{s}}):=\mathfrak{S}_{u}({\textbf{t}},{\textbf{s}})\mathfrak{S}_{w_{0}w}({\textbf{t}},{\textbf{s}}^{\prime})$ is the double Richardson polynomial, where ${\textbf{s}}^{\prime}=(s_{n},\ldots,s_{1})$ denotes the reverse of ${\textbf{s}}=(s_{1},\ldots,s_{n})$.
We point out that the matrix Richardson variety $D_{u}^{w}$ is a reduced and irreducible $T$-subvariety of ${\rm Mat}_{n,n}$. We have that $D_{u}^{w}$ is nonempty if and only if $w\geq u$ in the Bruhat order, and when it is nonempty, it has dimension $\dim(D_{u}^{w})=\ell(w)-\ell(u)$. For details on Richardson varieties, the reader is referred to, e.g., [Ric92, Bri05, Spe23].
Our main result in this section is the following theorem.
Theorem 5.2.
For two permutations $u,w\in S_{n}$ with $w\geq u$ in Bruhat order, the following statements hold:
  1. (i)
    The double Richardson polynomial $\mathfrak{R}_{w/u}({\textbf{t}},{\textbf{s}})$ represents the equivariant class of the matrix Richardson variety $D_{u}^{w}$ in $\text{H}_{T}^{\bullet}({\rm Mat}_{n,n})$.
  2. (ii)
    The (sign-changed) double Richardson polynomial $\mathfrak{R}_{w/u}({\textbf{t}},-{\textbf{s}})$ is a covolume polynomial.
Proof.
(i) We have the equality $[D_{u}^{w}]^{T}=[D_{u}\cap D^{w}]^{T}=[D_{u}]^{T}\cdot[D^{w}]^{T}$ in $\text{H}_{T}^{\bullet}({\rm Mat}_{n,n})$. Therefore Section missing 5 and Theorem 5.1 imply that $\mathfrak{S}_{u}({\textbf{t}},{\textbf{s}})\mathfrak{S}_{w_{0}w}({\textbf{t}},{\textbf{s}}^{\prime})=\mathfrak{R}_{w/u}({\textbf{t}},{\textbf{s}})$ represents the equivariant class of the matrix Richardson variety $D_{u}^{w}$.
(ii) By utilizing part (i) and Section missing 4, it follows that $\mathfrak{R}_{w/u}({\textbf{t}},-{\textbf{s}})$ is a covolume polynomial. ∎
The double Richardson polynomial specializes to many polynomials of interest:
  1. (i)
    $\mathfrak{R}_{w_{0}/u}({\textbf{t}},{\textbf{s}})=\mathfrak{S}_{u}({\textbf{t}},{\textbf{s}})$ is the double Schubert polynomial.
  2. (ii)
    $\mathfrak{R}_{w_{0}/u}({\textbf{t}},\mathbf{0})=\mathfrak{S}_{u}({\textbf{t}},\mathbf{0})=\mathfrak{S}_{u}({\textbf{t}})$ is the ordinary Schubert polynomial.
  3. (iii)
    We say that $\mathfrak{R}_{w/u}({\textbf{t}})=\mathfrak{R}_{w/u}({\textbf{t}},\mathbf{0})$ is the (ordinary) Richardson polynomial.
The next proposition shows that dually Lorentzian polynomials are discretely log-concave.
[…]
If $h=\sum_{\mathbf{n}}a_{\mathbf{n}}{\textbf{t}}^{\mathbf{n}}$ is a dually Lorentzian polynomial, we have that
$a_{\mathbf{n}}^{2}\;\,\geq\;\,a_{{\mathbf{n}}+{\mathbf{e}}_{i}-{\mathbf{e}}_{j}}a_{{\mathbf{n}}-{\mathbf{e}}_{i}+{\mathbf{e}}_{j}}.$
Proof.
Since $N\left({\textbf{t}}^{\mathbf{m}}h(t_{1}^{-1},\ldots,t_{p}^{-1})\right)=\sum_{{\mathbf{n}}}\frac{a_{\mathbf{n}}}{({\mathbf{m}}-{\mathbf{n}})!}{\textbf{t}}^{{\mathbf{m}}-{\mathbf{n}}}$ is Lorentzian for some ${\mathbf{m}}\in\mathbb{N}^{p}$ large enough, by [BH20, Proposition 4.4] we get the inequality in the statement. ∎
The following remark states that the truncation of a dually Lorentzian polynomial is dually Lorentzian. This result also follows from [RSW23, §3], but here we give a short self-contained proof.
[…]
Let $P({\textbf{t}})=\sum_{\mathbf{n}}a_{\mathbf{n}}{\textbf{t}}^{{\mathbf{n}}}$ be a dually Lorentzian polynomial. Let ${\mathbf{w}}=(w_{1},\ldots,w_{n})\in\mathbb{N}^{p}$. Then the truncation
$P^{\prime}({\textbf{t}})\;=\;\sum_{{\mathbf{n}}\leq{\mathbf{w}}}a_{\mathbf{n}}{\textbf{t}}^{{\mathbf{n}}}$
is dually Lorentzian.
Proof.
Take ${\mathbf{m}}\in\mathbb{N}^{p}$ large enough such that ${\mathbf{w}}\leq{\mathbf{m}}$ and
$Q({\textbf{t}})\;=\;N\left({\textbf{t}}^{\mathbf{m}}P(t_{1}^{-1},\ldots,t_{p}^{-1})\right)\;=\;\sum_{{\mathbf{n}}}\frac{a_{\mathbf{n}}}{({\mathbf{m}}-{\mathbf{n}})!}{\textbf{t}}^{{\mathbf{m}}-{\mathbf{n}}}$
is a Lorentzian polynomial. Then
$\partial_{t_{1}}^{m_{1}-w_{1}}\cdots\partial_{t_{p}}^{m_{p}-w_{p}}Q({\textbf{t}})\;=\;\sum_{{\mathbf{n}}\leq{\mathbf{w}}}\frac{a_{\mathbf{n}}}{({\mathbf{w}}-{\mathbf{n}})!}{\textbf{t}}^{{\mathbf{w}}-{\mathbf{n}}}$
is also Lorentzian, and consequently
$\sum_{{\mathbf{n}}\leq{\mathbf{w}}}a_{\mathbf{n}}{\textbf{t}}^{{\mathbf{n}}+{\mathbf{m}}-{\mathbf{w}}}\;=\;{\textbf{t}}^{{\mathbf{m}}-{\mathbf{w}}}P^{\prime}({\textbf{t}})$
is dually Lorentzian. Finally, §missing 2.3 implies that $P^{\prime}({\textbf{t}})$ is dually Lorentzian. ∎
We now obtain some consequences for certain polynomials of interest in algebraic combinatorics.
[…]
The following polynomials have M-convex support and are discretely log-concave:
  1. (i)
    (sign-changed) Double Richardson polynomials $\mathfrak{R}_{w/u}({\textbf{t}},-{\textbf{s}})$.
  2. (ii)
    Richardson polynomials $\mathfrak{R}_{w/u}({\textbf{t}})$.
  3. (iii)
    (sign-changed) Double Schubert polynomials $\mathfrak{S}_{u}({\textbf{t}},-{\textbf{s}})$.
  4. (iv)
    Schubert polynomials $\mathfrak{S}_{u}({\textbf{t}})$.
  5. (v)
    Any truncation of the above polynomials.
Proof.
By Theorem 5.2 and Section missing 5, these polynomials and their truncations are dually Lorentzian.
Since all these polynomials are dually Lorentzian and Lorentzian polynomials have M-convex support, [Sch03, §44.6f] implies that the support of all these polynomials is M-convex (also, see [Alu24, Corollary 2.12]). The discrete log-concavity statement follows from Section missing 5. ∎
We now compare the Richardson polynomial $\mathfrak{R}_{w/u}({\textbf{t}})$ with the skew Schubert polynomial of Lenart and Sottile [LS03] (also, see [WY14, Remark 2.16]).
[…]
Let $w\geq u$ be two permutations in $S_{n}$. Lenart and Sottile [LS03] defined the skew Schubert polynomial $\mathfrak{S}_{w/u}$ as the polynomial representative in normal form of the class of the Richardson variety $X_{u}^{w}=X_{u}\cap X^{w}$ in the cohomology ring $\text{H}^{\bullet}({{\mathcal{F}\ell}_{n}})$ of the full flag variety ${{\mathcal{F}\ell}_{n}}$. The Borel presentation
$\text{H}^{\bullet}({{\mathcal{F}\ell}_{n}})\;=\;\frac{{\mathbb{Z}}[t_{1},\ldots,t_{n}]}{\left(e_{1}(t_{1},\ldots,t_{n}),\ldots,e_{n}(t_{1},\ldots,t_{n})\right)}$
(where $e_{i}(t_{1},\ldots,t_{n})$ is the $i$-th elementary symmetric polynomial) yields a ${\mathbb{Z}}$-basis given by the monomials $t_{1}^{a_{1}}\cdots t_{n}^{a_{n}}$ with $a_{i}\leq n-i$. We say that a polynomial representative is in normal form if it can be written as a ${\mathbb{Z}}$-linear combination of this distinguished monomial basis. Let $n=4$ and consider the permutations $w=3412>2143=u$. The Richardson polynomial is given by
$\mathfrak{R}_{3412/2143}({\textbf{t}})\;=\;\mathfrak{S}_{2143}({\textbf{t}})^{2}\;=\;t_{1}^{4}+2\,t_{1}^{3}t_{2}+t_{1}^{2}t_{2}^{2}+2\,t_{1}^{3}t_{3}+2\,t_{1}^{2}t_{2}t_{3}+t_{1}^{2}t_{3}^{2}.$
However, the skew Schubert polynomial (the polynomial representative in normal form) is given by
$\mathfrak{S}_{3412/2143}({\textbf{t}})\;=\;t_{1}^{3}t_{2}\,+\,t_{1}^{3}t_{3}\,+\,t_{1}^{2}t_{2}t_{3}.$
Due to Section missing 5, if the skew Schubert polynomial $\mathfrak{S}_{w/u}({\textbf{t}})$ is a truncation of the Richardson polynomial $\mathfrak{R}_{w/u}({\textbf{t}})$, then $\mathfrak{S}_{w/u}({\textbf{t}})$ is a dually Lorentzian polynomial. However, as in Section missing 5, $\mathfrak{S}_{w/u}({\textbf{t}})$ may not coincide with a truncation of $\mathfrak{R}_{w/u}({\textbf{t}})$. Thus we should ask the question below.
[…]
Let $w\geq u$ be two permutations in $S_{n}$. We ask the following:
  1. (a)
    Is the support of $\mathfrak{S}_{w/u}({\textbf{t}})$ an M-convex set?
  2. (b)
    Is $\mathfrak{S}_{w/u}({\textbf{t}})$ discretely log-concave?
  3. (c)
    Is $\mathfrak{S}_{w/u}({\textbf{t}})$ a dually Lorentzian polynomial?
  4. (d)
    Is $\mathfrak{S}_{w/u}({\textbf{t}})$ a Lorentzian polynomial?
6. Macaulay dual generators of cohomology rings
In this section, we study the Macaulay dual generators of cohomology rings of smooth complex algebraic varieties. We prove that under certain positivity assumptions the Macaulay dual generators are denormalized Lorentzian polynomials.
6.1. Gorenstein algebras over a base ring
Before presenting our results on cohomology rings, which is our main interest, we develop basic ideas regarding the notion of Gorenstein algebras over a base ring. (These developments are probably known to the experts but we could not find a suitable reference for us.) For details on Gorenstein rings and duality results, the reader is referred to [Eis95, Chapter 21], [BH98, Chapter 3]. Throughout this subsection we use the following setup.
[…]
Let $A$ be a Noetherian ring (always assumed to be commutative and with identity) and $R$ be a positively graded finitely generated algebra over $R_{0}=A$. Choose a positively graded polynomial ring $S=A[x_{1},\ldots,x_{n}]$ with a graded surjection $S\twoheadrightarrow R$ and write $R\cong S/I$ for some homogeneous ideal $I\subset S$. Let ${\mathfrak{m}}:=(x_{1},\ldots,x_{n})\subset S$ be the graded irrelevant ideal. Let $\delta_{i}:=\deg(x_{i})>0$ and $\delta:=\delta_{1}+\cdots+\delta_{n}$. For any $\mathfrak{p}\in{\text{Spec}}(A)$, we denote by $\kappa(\mathfrak{p}):=A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ the residue field at $\mathfrak{p}$ and we set $S(\mathfrak{p}):=S\otimes_{A}\kappa(\mathfrak{p})\cong\kappa(\mathfrak{p})[x_{1},\ldots,x_{n}]$ and $R(\mathfrak{p}):=R\otimes_{A}\kappa(\mathfrak{p})$.
We are interested in the following relative notion of Gorenstein.
[…]
We say that $R$ is a Gorenstein algebra over $A$ if the natural morphism $f\colon{\text{Spec}}(R)\rightarrow{\text{Spec}}(A)$ is a Gorenstein morphism (see [Sta24, Tag 0C02]). This means that $R$ is $A$-flat and the fiber $R(\mathfrak{p})=R\otimes_{A}\kappa(\mathfrak{p})$ is a Gorenstein ring for all $\mathfrak{p}\in{\text{Spec}}(A)$. We also say that $R$ is an Artinian Gorenstein algebra over $A$ if $f\colon{\text{Spec}}(R)\rightarrow{\text{Spec}}(A)$ is a finite Gorenstein morphism.
For a graded $S$-module $M$, we denote the $A$-relative graded Matlis dual by
$M^{\vee_{A}}\;=\;{}^{*}{\text{Hom}}_{A}(M,A)\;:=\;\bigoplus_{\nu\in{\mathbb{Z}}}{\text{Hom}}_{A}\left({M}_{-\nu},A\right).$
[…]
Let $L,M,N$ be finitely generated graded $R$-modules that are $A$-flat. Then:
  1. (i)
    For all $\nu\in{\mathbb{Z}}$, the graded part $M_{\nu}$ is a finitely generated flat $A$-module, and thus $A$-projective.
  2. (ii)
    $\left(M^{\vee_{A}}\right)^{\vee_{A}}\cong M$.
  3. (iii)
    If $0\rightarrow L\rightarrow M\rightarrow N\rightarrow 0$ is a short exact sequence, then $0\rightarrow N^{\vee_{A}}\rightarrow M^{\vee_{A}}\rightarrow L^{\vee_{A}}\rightarrow 0$ is also exact.
[…]
All the fibers $R(\mathfrak{p})=R\otimes_{A}\kappa(\mathfrak{p})$ of $f\colon{\text{Spec}}(R)\rightarrow{\text{Spec}}(A)$ have the same dimension in either of the following two cases:
  1. (i)
    $f$ is a finite morphism.
  2. (ii)
    $f$ is flat and ${\text{Spec}}(A)$ is connected (e.g., $A$ is a domain).
Proof.
(i) In this case it is clear that all the fibers $R(\mathfrak{p})$ have Krull dimension 00.
(ii) Notice that each graded part $R_{\nu}$ is a locally-free $A$-module of constant rank (see, e.g., [Sta24, Tag 00NX]). Thus the Hilbert function of the graded $\kappa(\mathfrak{p})$-algebra $R(\mathfrak{p})$ is the same for all $\mathfrak{p}\in{\text{Spec}}(A)$. ∎
[…]
Let $M$ be a finitely generated $R$-module. Then $M=0$ if and only if $M\otimes_{A}\kappa(\mathfrak{p})=0$ for all $\mathfrak{p}\in{\text{Spec}}(A)$.
[…]
A Cohen–Macaulay ring $S$ with a canonical module $\omega_{S}$ is Gorenstein if and only if $\omega_{S}$ is generated locally by one element (i.e., $\omega_{S_{P}}\cong\omega_{S}\otimes_{S}S_{P}$ is generated by one element for all $P\in{\text{Spec}}(S)$).
The next proposition is similar to [CR23, Lemma 2.10].
[…]
Assume that $f\colon{\text{Spec}}(R)\rightarrow{\text{Spec}}(A)$ is flat and that either ${\text{Spec}}(A)$ is connected or $f$ is finite. Let $e$ be the common dimension of the fibers of $f$ (see §missing 6.1). Then $R$ is Gorenstein over $A$ if and only if the following two conditions are satisfied:
  1. (a)
    $\text{Ext}_{S}^{i}(R,S)=0$ for all $0\leq i\leq n$ such that $i\neq n-e$.
  2. (b)
    $\text{Ext}_{S}^{n-e}(R,S)$ is $A$-flat and it is generated locally by one element as an $R$-module.
Proof.
Here we use the property of exchange for local Ext modules (see [LK79, Theorem A.5], [AK80, Theorem 1.9]). For any $\mathfrak{p}\in{\text{Spec}}(A)$, consider the natural base change maps
$\gamma_{\mathfrak{p}}^{i}\colon\;\text{Ext}_{S}^{i}(R,S)\otimes_{A}\kappa(\mathfrak{p})\;\rightarrow\;\text{Ext}_{S(\mathfrak{p})}^{i}(R(\mathfrak{p}),S(\mathfrak{p})).$
We have the following two important properties:
  1. (i)
    If $\gamma_{\mathfrak{p}}^{i}$ is surjective, then it is an isomorphism.
  2. (ii)
    If $\gamma_{\mathfrak{p}}^{i}$ is surjective, then $\gamma_{\mathfrak{p}}^{i-1}$ is surjective if and only if $\text{Ext}_{S}^{i}(R,S)\otimes_{A}A_{\mathfrak{p}}$ is $A_{\mathfrak{p}}$-flat.
For all $i>n$, since each $\gamma_{\mathfrak{p}}^{i}$ is surjective by Hilbert’s syzygy theorem, we obtain the vanishing $\text{Ext}_{S}^{i}(R,S)=0$ (see §missing 6.1). For all $\mathfrak{p}\in{\text{Spec}}(A)$, recall that the fiber $R(\mathfrak{p})$ is a Cohen–Macaulay ring if and only if $\text{Ext}_{S(\mathfrak{p})}^{i}(R(p),S(\mathfrak{p}))=0$ for all $i\neq n-e$ (this follows for instance by the local duality theorem).
Let $i>n-e$. If each $\gamma_{\mathfrak{p}}^{i+1}$ is surjective and $\text{Ext}_{S}^{i+1}(R,S)=0$, it follows that $\gamma_{\mathfrak{p}}^{i}$ is surjective, and this implies that $\text{Ext}_{S}^{i}(R,S)=0$ if and only if $\text{Ext}_{S(\mathfrak{p})}^{i}(R(\mathfrak{p}),S(\mathfrak{p}))=0$ for all $\mathfrak{p}\in{\text{Spec}}(A)$. Hence, by descending induction on $i$, we obtain $\text{Ext}_{S}^{i}(R,S)=0$ for all $i>n-e$ if and only if $\text{Ext}_{S(\mathfrak{p})}^{i}(R(\mathfrak{p}),S(\mathfrak{p}))=0$ for all $i>n-e$ and $\mathfrak{p}\in{\text{Spec}}(A)$.
So we assume that $\text{Ext}_{S}^{i}(R,S)=0$ for all $i>n-e$ (and, equivalently, that $\text{Ext}_{S(\mathfrak{p})}^{i}(R(\mathfrak{p}),S(\mathfrak{p}))=0$ for all $i>n-e$ and $\mathfrak{p}\in{\text{Spec}}(A)$).
Since each $\gamma_{\mathfrak{p}}^{n-e+1}$ is surjective and $\text{Ext}_{S}^{n-e+1}(R,S)=0$, it follows that each $\gamma_{\mathfrak{p}}^{n-e}$ is surjective. In turn, this implies that $\gamma_{\mathfrak{p}}^{n-e-1}$ is surjective for all $\mathfrak{p}\in{\text{Spec}}(A)$ if and only if $\text{Ext}_{S}^{n-e}(R,S)$ is $A$-flat. Similarly, by descending induction on $i$, we can show that $\text{Ext}_{S(\mathfrak{p})}^{i}(R(\mathfrak{p}),S(\mathfrak{p}))=0$ for all $i<n-e$ and $\mathfrak{p}\in{\text{Spec}}(A)$ if and only if $\text{Ext}_{S}^{i}(R,S)=0$ for all $i<n-e$ and $\text{Ext}_{S}^{n-e}(R,S)$ is $A$-flat.
Therefore, in our current setting, we have shown that $f\colon{\text{Spec}}(R)\rightarrow{\text{Spec}}(A)$ is a Cohen–Macaulay morphism (see [Sta24, Tag 045Q]) if and only if $\text{Ext}_{S}^{i}(R,S)=0$ for all $0\leq i\leq n$ with $i\neq n-e$ and $\text{Ext}_{S}^{n-e}(R,S)$ is $A$-flat.
Let $M=\text{Ext}_{S}^{n-e}(R,S)$. To conclude the proof, we may assume that $f$ is a Cohen–Macaulay morphism, and we need to show that $M$ can be generated locally by one element if and only if $M\otimes_{A}\kappa(\mathfrak{p})\cong\text{Ext}_{S(\mathfrak{p})}^{n-e}(R(\mathfrak{p}),S(\mathfrak{p}))$ can be generated by one element for all $\mathfrak{p}\in{\text{Spec}}(A)$. This can be proved using basic properties of Fitting ideals (see [Eis95, §20.2]). Indeed, the claim follows by [Eis95, Proposition 20.6] because ${\rm Fitt}_{1}(M)=R$ if and only if ${\rm Fitt}_{1}(M\otimes_{A}\kappa(\mathfrak{p}))=\left({\rm Fitt}_{1}(M)\right)R(\mathfrak{p})=R(\mathfrak{p})$ for all $\mathfrak{p}\in{\text{Spec}}(A)$ (again, see §missing 6.1). So the proof is complete. ∎
Motivated by the above proposition, when $R$ is Gorenstein over $A$ and $e$ is the common dimension of the fibers, we say that
$\omega_{R/A}\;:=\;\text{Ext}_{S}^{n-e}\big{(}R,S(-\delta)\big{)}$
is a relative canonical module of $R$ over $A$.
Assume that $R$ is a finitely generated $A$-module. Let $d>0$ be a positive integer and $\rho\colon R\rightarrow A(-d)$ be a graded $A$-linear map (following tradition we call this map an orientation). We say that $R$ satisfies Poincaré duality with respect to the orientation $\rho$ if
$R\otimes_{A}R\;\rightarrow\;A(-d),\quad r_{1}\otimes r_{2}\mapsto\rho(r_{1}r_{2})$
is a perfect pairing. This means that the pairing induces a graded isomorphism $R\cong{}^{*}{\text{Hom}}_{A}(R,A)(-d)$ of $R$-modules. In particular, we have perfect pairings $R_{i}\otimes_{A}R_{d-i}\rightarrow A$ for all $0\leq i\leq d$. We also denote by
$\rho\colon R_{d}\xrightarrow{\;\cong\;}A$
the induced isomorphism. Since we are assuming that $R$ is a finitely generated $A$-module, we actually have that ${}^{*}{\text{Hom}}_{A}(R,A)$ coincides with ${\text{Hom}}_{A}(R,A)$ (see [BH98, Exercise 1.5.19]), but we prefer the notation ${}^{*}{\text{Hom}}_{A}(R,A)$ to stress our interest in the graded parts of $R$. We consider the inverse polynomial ring $T=A[y_{1},\ldots,y_{n}]$ where $y_{i}$ is identified with $x_{i}^{-1}$. The $S$-module structure of $T$ is given by setting that
$x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}\,\cdot\,y_{1}^{\beta_{1}}\cdots y_{n}^{\beta_{n}}\;\;=\;\;\begin{cases}y_{1}^{\beta_{1}-\alpha_{1}}\cdots y_{n}^{\beta_{n}-\alpha_{n}}&\text{if $\beta_{i}\geq\alpha_{i}$ for all $1\leq i\leq n$}\\ 0&\text{otherwise.}\end{cases}$
Then $T$ is a negatively graded polynomial ring with $\deg(y_{i})=-\delta_{i}$. We have the natural isomorphisms
$T\;\cong\;{}^{*}{\text{Hom}}_{A}(S,A)\;\cong\;\text{H}_{{\mathfrak{m}}}^{n}\big{(}S\big{)}(-\delta)$
of graded $S$-modules. Recall the basic computation $\text{H}_{{\mathfrak{m}}}^{n}(S)\cong\frac{1}{x_{1}\ldots x_{n}}A[x_{1}^{-1},\ldots,x_{n}^{-1}]$.
The following theorem extends well-known results about Artinian Gorenstein algebras (over a field) to our current relative setting over a Noetherian base ring.
Theorem 6.1.
Assume that $f\colon{\text{Spec}}(R)\rightarrow{\text{Spec}}(A)$ is a finite flat morphism and that $R$ satisfies Poincaré duality with respect to an orientation $\rho\colon R\rightarrow A(-d)$. Then the following statements hold:
  1. (i)
    $R$ is Artinian Gorenstein over $A$.
  2. (ii)
    We have the isomorphisms $\omega_{R/A}=\text{Ext}_{S}^{n}(R,S(-\delta))\cong{}^{*}{\text{Hom}}_{A}(R,A)\cong R(d)$.
  3. (iii)
    Consider the inverse polynomial
    $G_{R}(y_{1},\ldots,y_{n})\;=\;\sum_{\alpha_{1}\delta_{1}+\cdots+\alpha_{n}\delta_{n}=d}\rho\big{(}x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}\big{)}y_{1}^{\alpha_{1}}\cdots y_{n}^{\alpha_{n}}\;\;\in\;\;T=A[y_{1},\ldots,y_{n}].$
    Then the presenting ideal of $R=S/I$ is given as the annihilator of $G_{R}$:
    $I\;=\;\{g\in S\mid g\cdot G_{R}=0\}.$
Proof.
By the Hom-tensor adjunction, we have the isomorphism
${\text{Hom}}_{R}\left(R,{}^{*}{\text{Hom}}_{A}(R,A(-d))\right)\;\cong\;{}^{*}{\text{Hom}}_{A}\left(R\otimes_{A}R,A(-d)\right).$
Thus having a perfect pairing $R\otimes_{A}R\rightarrow A(-d)$ is the same as having $R\cong{}^{*}{\text{Hom}}_{A}(R,A)(-d)$.
Since $R=\text{H}_{{\mathfrak{m}}}^{0}(R)$ is $A$-flat and $\text{H}_{{\mathfrak{m}}}^{i}(R)=0$ for $i\geq 1$, the generic version of the local duality theorem given in [CR23, Theorem A] yields the isomorphisms $\text{Ext}_{S}^{n-i}(R,S(-\delta))\cong{}^{*}{\text{Hom}}_{A}(\text{H}_{{\mathfrak{m}}}^{i}(R),A)$ for all $i$. Therefore, $R$ is Gorenstein over $A$ by §missing 6.1 and we have the isomorphisms $\omega_{R/A}=\text{Ext}_{S}^{n}(R,S(-\delta))\cong{}^{*}{\text{Hom}}_{A}(R,A)\cong R(d)$. So the proofs of part (i) and (ii) are complete.
Let $E=I^{\perp_{A}}=\{G\in T\mid g\cdot G=0\text{ for all $g\in I$}\}$ be the (relative) inverse system of $I$. We have the natural isomorphisms
$E\;\cong\;{\text{Hom}}_{S}\big{(}R,T\big{)}\;\cong\;{\text{Hom}}_{S}\big{(}R,{}^{*}{\text{Hom}}_{A}(S,A)\big{)}\;\cong\;{}^{*}{\text{Hom}}_{A}(R,A)=R^{\vee_{A}}.$
The isomorphism $E\cong R^{\vee_{A}}$ can be obtained by dualizing the short exact sequence $0\rightarrow I\rightarrow S\rightarrow R\rightarrow 0$ and writing
[…]
(see §missing 6.1). By part (ii), there is an inverse polynomial $G\in T=A[y_{1},\ldots,y_{n}]$ of degree $-d$ that generates $E$ as an $R$-module. Then the dual short exact sequence $0\rightarrow I\rightarrow S\rightarrow E^{\vee_{A}}\rightarrow 0$ shows that
$I\;=\;\{g\in S\mid g\cdot G=0\}.$
To conclude the proof, it suffices to show that $G_{R}$ generates the $A$-module $E_{-d}$. By construction, for any homogeneous polynomial $g\in S_{d}$, we have that
$g\cdot G_{R}\;=\;\rho\left(\overline{g}\right),$
where $\overline{g}$ denotes the class of $g\in S_{d}$ in $R_{d}=S_{d}/I_{d}$. This implies that under the isomorphism $E_{-d}\cong{\text{Hom}}_{A}(R_{d},A)$, the inverse polynomial $G_{R}\in E_{-d}$ corresponds to the isomorphism $\rho\in{\text{Hom}}_{A}(R_{d},A)$. On the other hand, under the isomorphism ${\text{Hom}}_{A}(R_{d},A)\cong A$, the isomorphism $\rho$ should correspond to a unit in $A$. This shows that $G_{R}$ is a generator of $E_{-d}$, and so the proof of the theorem is complete. ∎
[…]
Following standard notation, we say that the inverse polynomial $G_{R}\in T=A[y_{1},\ldots,y_{n}]$ presented in Theorem 6.1 is the Macaulay dual generator of $R$ over $A$.
6.2. Cohomology rings as Artinian Gorenstein algebras over ${\mathbb{Z}}$
In this subsection, we apply our developments in §missing 6.1 to cohomology rings. Our main result in this direction is the following theorem.
Theorem 6.2.
Let $X$ be a $d$-dimensional smooth complex algebraic variety. Suppose that the cohomology ring $R=\bigoplus_{i=0}^{d}\text{H}^{2i}(X,{\mathbb{Z}})$ is a flat ${\mathbb{Z}}$-algebra (i.e., it is ${\mathbb{Z}}$-torsion-free). Let $\rho\colon R_{d}=\text{H}^{2d}(X,{\mathbb{Z}})\rightarrow{\mathbb{Z}}$ be the natural degree map. Choose a graded presentation $R\cong S/I$ where $S={\mathbb{Z}}[x_{1},\ldots,x_{n}]$, $\delta_{i}=\deg(x_{i})>0$, and $I\subset S$ is a homogeneous ideal. Let $\delta=\delta_{1}+\cdots+\delta_{n}$. Then the following statements hold:
  1. (i)
    $R$ is Artinian Gorenstein over ${\mathbb{Z}}$.
  2. (ii)
    We have the isomorphisms $\omega_{R/{\mathbb{Z}}}=\text{Ext}_{S}^{n}(R,S(-\delta))\cong{}^{*}{\text{Hom}}_{\mathbb{Z}}(R,{\mathbb{Z}})\cong R(d)$.
  3. (iii)
    Consider the inverse polynomial ring $T={\mathbb{Z}}[y_{1},\ldots,y_{n}]$ with the identification $y_{i}=x_{i}^{-1}$. The ideal $I\subset S$ is given as the annihilator
    $I\;=\;\{g\in S\mid g\cdot G_{R}=0\}$
    of the inverse polynomial
    $G_{R}(y_{1},\ldots,y_{n})\;=\;\sum_{\alpha_{1}\delta_{1}+\cdots+\alpha_{n}\delta_{n}=d}\rho\big{(}x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}\big{)}y_{1}^{\alpha_{1}}\cdots y_{n}^{\alpha_{n}}\;\;\in\;\;T={\mathbb{Z}}[y_{1},\ldots,y_{n}].$
  4. (iv)
    Assume that $X$ is complete and that each $x_{i}$ is equal to the first Chern class $c_{1}(L_{i})$ of a nef line bundle $L_{i}$ on $X$. Then the normalization
    $N(G_{R})\;\in\;\mathbb{R}[y_{1},\ldots,y_{n}]$
    of $G_{R}$ is a Lorentzian polynomial.
Proof.
Parts (i), (ii), and (iii) follow from Theorem 6.1 because $\text{H}^{\bullet}(X,{\mathbb{Z}})$ satisfies Poincaré duality and we are assuming it is ${\mathbb{Z}}$-torsion-free.
We now concentrate on proving part (iv). Notice that the normalization
$N(G_{R})\;=\;\frac{1}{d!}\int\big{(}x_{1}y_{1}+\cdots+x_{n}y_{n}\big{)}^{d}\cap[X]\;\;\in\;\;\mathbb{R}[y_{1},\ldots,y_{n}]$
of $G_{R}$ is the volume polynomial of $x_{1},\ldots,x_{n}$ divided by $d!$. By using Chow’s lemma, we can find a proper birational surjective morphism $f\colon X^{\prime}\rightarrow X$ where $X^{\prime}$ is a projective variety. Let $L_{i}^{\prime}=f^{*}(L_{i})$ and $x_{i}^{\prime}=c_{1}(L_{i}^{\prime})$. The projection formula yields the equality
$N(G_{R})\;=\;\frac{1}{d!}\int\big{(}x_{1}^{\prime}y_{1}+\cdots+x_{n}^{\prime}y_{n}\big{)}^{d}\cap[X^{\prime}]\;\;\in\;\;\mathbb{R}[y_{1},\ldots,y_{n}].$
Since each $L_{i}^{\prime}$ is nef on $X^{\prime}$, we have that $N(G_{R})$ is a Lorentzian polynomial due to [BH20, Theorem 4.6]. ∎
We close this subsection with some specific computations.
[…]
By considering the Borel presentation
$R\;=\;\text{H}^{\bullet}(\mathcal{F}\ell_{3},{\mathbb{Z}})\;=\;\frac{{\mathbb{Z}}[x_{1},x_{2},x_{3}]}{\left(x_{1}+x_{2}+x_{3},\,x_{1}x_{2}+x_{1}x_{3}+x_{2}x_{3},\,x_{1}x_{2}x_{3}\right)}$
of the flag variety $\mathcal{F}\ell_{3}$, we obtain the Macaulay dual generator
$G_{R}(y_{1},y_{2},y_{3})\;=\;-y_{1}^{2}y_{2}+y_{1}^{2}y_{3}+y_{1}y_{2}^{2}-y_{1}y_{3}^{2}-y_{2}^{2}y_{3}+y_{2}y_{3}^{2}.$
We know that the elements $\widehat{x}_{1}=x_{1}$, $\widehat{x}_{2}=x_{1}+x_{2}$ and $\widehat{x}_{3}=x_{1}+x_{2}+x_{3}$ correspond to nef line bundles on $\mathcal{F}\ell_{3}$ (see, e.g., [Ful97, Chapter 10]). By considering the following alternative presentation
$R\;=\;\text{H}^{\bullet}(\mathcal{F}\ell_{3},{\mathbb{Z}})\;=\;\frac{{\mathbb{Z}}[\widehat{x}_{1},\widehat{x}_{2},\widehat{x}_{3}]}{\left(\widehat{x}_{3},\,\;-\widehat{x}_{1}^{2}+\widehat{x}_{1}\widehat{x}_{2}-\widehat{x}_{2}^{2}+\widehat{x}_{2}\widehat{x}_{3},\,\;\widehat{x}_{1}^{2}\widehat{x}_{2}-\widehat{x}_{1}\widehat{x}_{2}^{2}-\widehat{x}_{1}^{2}\widehat{x}_{3}+\widehat{x}_{1}\widehat{x}_{2}\widehat{x}_{3}\right)},$
we obtain the following Macaulay dual generator
$G_{R}(\widehat{y}_{1},\widehat{y}_{2},\widehat{y}_{3})\;=\;\widehat{y}_{1}^{2}\widehat{y}_{2}+\widehat{y}_{1}\widehat{y}_{2}^{2}$
whose normalization is a Lorentzian polynomial (as predicted by Theorem 6.2(iv)).
[…]
The cohomology ring of the Grassmannian ${\rm Gr}(2,4)$ is given by
$R\;=\;\text{H}^{\bullet}({\rm Gr}(2,4),{\mathbb{Z}})\;=\;\frac{{\mathbb{Z}}[x_{1},x_{2}]}{\left(x_{1}^{3}-2x_{1}x_{2},\,\,x_{1}^{2}x_{2}-x_{2}^{2}\right)},$
where $x_{i}$ is the $i$-th Chern class $c_{i}(\mathcal{S})$ of the universal subbundle $\mathcal{S}$ on ${\rm Gr}(2,4)$. Then the Macaulay dual generator is given by
[…]
Finally, we provide a characteristic-free (over ${\mathbb{Z}}$) extension of the celebrated result of Khovanskii and Pukhlikov [KP92] showing that the cohomology ring (over $\mathbb{Q}$) of certain toric varieties can be expressed in terms of differential operators that annihilate the volume polynomial. We follow the notations in [CLS11].
We briefly recall the notion of mixed volume (for more details, the reader is referred to [Ewa96, Chapter IV], [CLO05, Chapter 7]). Let $P_{1},\dots,P_{n}$ be lattice polytopes in ${\mathbb{Z}}^{d}$ and ${\rm Vol}(P_{i})$ denote the Euclidean volume of $P_{i}$, where the unit hypercube has volume $1$. The volume of the Minkowski sum of polytopes ${\rm Vol}(y_{1}P_{1}+\cdots+y_{n}P_{n})$ is a homogeneous polynomial of degree $d$, and we write
${\rm Vol}(y_{1}P_{1}+\cdots+y_{n}P_{n})\;=\;\sum_{|\alpha|=d}\,\frac{d!}{\alpha!}\;\mathrm{MV}_{\alpha}(P_{1},\dots,P_{n})\,y_{1}^{\alpha_{1}}\cdots y_{n}^{\alpha_{n}}$
where $\mathrm{MV}_{\alpha}(P_{1},\dots,P_{n})$ is called the mixed volume of $(P_{1},\dots,P_{n})$ of type $\alpha$.
Let $X_{\Sigma}$ be a smooth complete toric variety and write $d=\dim(X_{\Sigma})$, $n=|\Sigma(1)|$, and $m={\rm rank}\left({\text{Pic}}(X_{\Sigma})\right)$. Let $\rho_{1},\dots,\rho_{n}$ be the one-dimensional rays in $\Sigma(1)$. Let $D_{i}$ be the divisor on $X_{\Sigma}$ associated to $\rho_{i}$. We have the following isomorphism
${\mathbb{Z}}[x_{1},\dots,x_{n}]/(\mathcal{I}+\mathcal{J})\;\;\xrightarrow{\;\cong\;}\;\;\text{H}^{\bullet}(X_{\Sigma},{\mathbb{Z}}),\qquad x_{i}\mapsto D_{i},$
where $\mathcal{I}=\left(x_{i_{1}},\dots,x_{i_{s}}\mid i_{j}\text{ are distinct and }\rho_{i_{1}}+\cdots+\rho_{i_{s}}\text{ is not a cone of }\Sigma\right)$ and $\mathcal{J}$ is the ideal generated by linear forms $\sum_{i=1}^{n}\langle m,u_{i}\rangle x_{i}$ for all $m\in M$ in the dual lattice $M$. Moreover, the cohomology ring $\text{H}^{\bullet}(X_{\Sigma},{\mathbb{Z}})$ is ${\mathbb{Z}}$-torsion-free. This explicit description of $\text{H}^{\bullet}(X_{\Sigma},{\mathbb{Z}})$ was obtained by Jurkiewicz [Jur80] and Danilov [Dan78]. The volume polynomial of $X_{\Sigma}$ is defined as
$V(y_{1},\dots,y_{n})\;:=\;\int_{X_{\Sigma}}\left(D_{1}y_{1}+\cdots+D_{n}y_{n}\right)^{d}.$
Assume that the classes of the divisors $\overline{D}_{1},\dots,\overline{D}_{m}$ on $X_{\Sigma}$ give a basis of ${\text{Pic}}(X_{\Sigma})$. Then the reduced volume polynomial of $X_{\Sigma}$ is given by
$\overline{V}(y_{1},\dots,y_{m})\;:=\;\int_{X_{\Sigma}}\left(\overline{D}_{1}y_{1}+\cdots+\overline{D}_{m}y_{m}\right)^{d}.$
Below we have our extension of the aforementioned result of Khovanskii and Pukhlikov [KP92].
[…]
Let $R=\bigoplus_{i=0}^{d}\text{H}^{2i}\left(X_{\Sigma},{\mathbb{Z}}\right)$. Under the Macaulay inverse system notation of Theorem 6.2, the following statements hold:
  1. (i)
    There are isomorphisms of ${\mathbb{Z}}$-algebras
    $R\;\cong\;{\mathbb{Z}}[x_{1},\dots,x_{n}]/I\;\cong\;{\mathbb{Z}}[x_{1},\dots,x_{m}]/J$
    where
    $I\;=\;\left\{g\in{\mathbb{Z}}[x_{1},\dots,x_{n}]\,\mid\,g\cdot N^{-1}\left(V(y_{1},\dots,y_{n})\right)=0\right\}\;=\;\mathcal{I}+\mathcal{J}$
    and
    $J\;=\;\left\{g\in{\mathbb{Z}}[x_{1},\dots,x_{m}]\,\mid\,g\cdot N^{-1}\left(\overline{V}(y_{1},\dots,y_{m})\right)=0\right\}.$
    Here $N^{-1}$ denotes the inverse of the normalization operator.
  2. (ii)
    If each $D_{i}$ is nef and $P_{i}=P_{D_{i}}$ is the polytope associated to $D_{i}$, then
    $G_{R}(y_{1},\ldots,y_{n})\;=\;\sum_{\alpha_{1}+\cdots+\alpha_{n}=d}\mathrm{MV}_{\alpha}(P_{1},\ldots,P_{n})\,y_{1}^{\alpha_{1}}\cdots y_{n}^{\alpha_{n}}.$
Proof.
(i) We concentrate on the isomorphism $R\cong{\mathbb{Z}}[x_{1},\ldots,x_{n}]/I$. From Theorem 6.2(iii), we have $I=\{g\in{\mathbb{Z}}[x_{1},\ldots,x_{n}]\mid g\cdot G_{R}=0\}$ and that the Macaulay dual generator is given by
$G_{R}(y_{1},\ldots,y_{n})\;=\;\sum_{\alpha_{1}+\cdots+\alpha_{n}=d}\left(\int_{X_{\Sigma}}D_{1}^{\alpha_{1}}\cdots D_{n}^{\alpha_{n}}\right)\,y_{1}^{\alpha_{1}}\cdots y_{n}^{\alpha_{n}}.$
Then the equality $I\;=\;\left\{g\in{\mathbb{Z}}[x_{1},\dots,x_{n}]\,\mid\,g\cdot N^{-1}\left(V(y_{1},\dots,y_{n})\right)=0\right\}$ follows by noticing that
${\rm Vol}(y_{1},\ldots,y_{n})\;=\;d!\,N\left(G_{R}(y_{1},\ldots,y_{n})\right).$
The isomorphism $R\cong{\mathbb{Z}}[x_{1},\ldots,x_{m}]/J$ follows verbatim.
(ii) From [CLS11, Theorem 13.4.3], we obtain the equality
$V(y_{1},\ldots,y_{n})\;=\;{\rm Vol}(y_{1}P_{1}+\cdots+y_{n}P_{n}).$
By comparing coefficients we get $\int_{X_{\Sigma}}D_{1}^{\alpha_{1}}\cdots D_{n}^{\alpha_{n}}=\mathrm{MV}_{\alpha}(P_{1},\ldots,P_{n})$. Therefore, the equality
$G_{R}(y_{1},\ldots,y_{n})\;=\;\sum_{\alpha_{1}+\cdots+\alpha_{n}=d}\mathrm{MV}_{\alpha}(P_{1},\ldots,P_{n})\,y_{1}^{\alpha_{1}}\cdots y_{n}^{\alpha_{n}}$
follows from Theorem 6.2(iii). ∎
[…]
Consider the Hirzebruch surface $\mathcal{H}_{r}$ and label the one dimensional rays in the fan as $\rho_{1}={\text{Cone}}(-{\mathbf{e}}_{1}+r{\mathbf{e}}_{2})$, $\rho_{2}={\text{Cone}}({\mathbf{e}}_{2})$, $\rho_{3}={\text{Cone}}({\mathbf{e}}_{1})$, and $\rho_{4}={\text{Cone}}(-{\mathbf{e}}_{2})$. The cohomology ring of $\mathcal{H}_{r}$ is given by
$R\;=\;\text{H}^{\bullet}(\mathcal{H}_{r},{\mathbb{Z}})\;\cong\;\frac{{\mathbb{Z}}[x_{1},x_{2},x_{3},x_{4}]}{(x_{1}x_{3},\,x_{2}x_{4},\,-x_{1}+x_{3},\,rx_{1}+x_{2}-x_{4})}.$
The Macaulay dual generator is given by
$G_{R}(y_{1},y_{2},y_{3},y_{4})\;=\;-ry_{2}^{2}+ry_{4}^{2}+y_{1}y_{2}+y_{2}y_{3}+y_{3}y_{4}+y_{1}y_{4}$
and the volume polynomial is given by
[…]
Alternatively, the cohomology ring of $\mathcal{H}_{r}$ can be represented as
$R\;=\;\text{H}^{\bullet}(\mathcal{H}_{r},{\mathbb{Z}})\;\cong\;\frac{{\mathbb{Z}}[x_{3},x_{4}]}{(x_{3}^{2},-rx_{3}x_{4}+x_{4}^{2})}.$
The Macaulay dual generator is given by
$G_{R}(y_{3},y_{4})\;=\;ry_{4}^{2}+y_{3}y_{4}$
and the reduced volume polynomial is given by
[…]
Since $D_{\rho_{3}}$ and $D_{\rho_{4}}$ are both nef divisors in $\mathcal{H}_{r}$ [CLS11, Example 6.3.23], we have that $\overline{V}(y_{3},y_{4})$ is a Lorentzian polynomial (as predicted by Theorem 6.2(iv)).